### v2

# Nilpotent structures of oriented neutral vector bundles

Abstract  In this paper, we study nilpotent structures of an oriented vector bundle $E$ of rank $4n$ with a neutral metric $h$ and an $h$-connection $\nabla$. We define $H$-nilpotent structures of $(E,h,\nabla)$ for a Lie subgroup $H$ of $SO(2n,2n)$ related to neutral hyperKähler structures. We observe that there exist a complex structure $I$ and paracomplex structures $J_{1}$, $J_{2}$ of $E$ such that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$ if and only if there exists an $H$-nilpotent structure of $(E,h,\nabla)$.
1 Introduction
Nilpotent structures of an oriented neutral vector bundle $(E,h)$ of rank $4n$ are analogues of complex structures and paracomplex structures. In particular, if $n=1$, then a nilpotent structure corresponds to a section of one of the light-like twistor spaces associated with $(E,h)$ (see [2], [3], [4] for the light-like twistor spaces) and a complex (respectively, paracomplex) structure preserving (respectively, reversing) the neutral metric $h$ corresponds to a section of one of the space-like (respectively, time-like) twistor spaces associated with $(E,h)$ (see [1], [5] for the space-like twistor spaces and see [1], [15], [16] for the time-like twistor spaces). A nilpotent structure $N$ gives a null structure on each fiber of $E$ and $h$ is null-Hermitian with respect to $N$ (see [11]).
A nilpotent structure $N$ gives a light-like subbundle $\pi_{N}$ of $(E,h)$ of rank $2n$. In addition, $N$ gives a section $\xi_{N}$ of $\bigwedge^{2n}\!E$, which is constructed by a local frame field of $\pi_{N}$ given by an admissible frame field of $N$. The fact that $\xi_{N}$ is well-defined is related to a Lie subgroup $G$ of $SO(2n,2n)$. The transition function between two admissible frame fields of $N$ is valued in $G$. It gives a local section of ${\rm End}\,\pi_{N}$ and its determinant is identically equal to one ([3]). If $n=1$, then the section of one of the light-like twistor spaces corresponding to $N$ is given by $(1/\sqrt{2})\xi_{N}$ ([2]).
Let $\nabla$ be an $h$-connection of $E$, i.e., a connection of $E$ satisfying $\nabla h=0$. Then the Walker condition of $N$ is defined, based on the definition of Walker manifolds (see [6], [8], [9], [11], [19] for Walker manifolds). The Walker condition of $N$ is characterized by ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ with a $1$-form $\alpha$ on $M$ for the connection $\hat{\nabla}$ of $\bigwedge^{2n}\!E$ induced by $\nabla$ ((a) in Proposition 3.1). Since $\nabla N=0$ means ^$\hat{\nabla}\xi_{N}=0$ ((b) in Proposition 3.1), a nilpotent structure $N$ parallel with respect to $\nabla$ satisfies the Walker condition (see [11]). If $n=1$, then $\nabla N=0$ is equivalent to ^$\hat{\nabla}\xi_{N}=0$ ([2]). There exists a nilpotent structure $N$ satisfying the Walker condition and ^$\hat{\nabla}\xi_{N}\not=0$ (Example 3.1, Example 3.2, Example 3.3). In addition, if $n\geq 2$, then there exists a nilpotent structure $N$ satisfying ^$\hat{\nabla}\xi_{N}=0$ and $\nabla N\not=0$ (Example 3.4). Such an example $N$ in Example 3.1 $\sim$ Example 3.4 admits a nilpotent structure $N^{\prime}$ of $(E,h)$ satisfying the Walker condition and $E=\pi_{N}\oplus\pi_{N^{\prime}}$.
In the present paper, we define special nilpotent structures related to a Lie subgroup $K$ of $SO(2n,2n)$. Each of them gives a principal $K$-bundle $P$ associated with $E$ and special admissible frame fields we use are given by local sections of $P$. In addition, $\nabla$ gives a connection in $P$, so that the connection form of $\nabla$ with respect to such a local section is valued in the Lie algebra of $K$. We call such a special nilpotent structure a $K$-nilpotent structure of $(E,h,\nabla)$. If $\nabla$ is flat and if $K\subset G$, then for a $K$-nilpotent structure of $(E,h,\nabla)$, we can find a local section $e$ of $P$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$ (Proposition 4.2). A $G$-nilpotent structure of $(E,h,\nabla)$ is just a nilpotent structure of $(E,h)$ parallel with respect to $\nabla$ (Proposition 4.3). Let $H$ be a Lie subgroup of $G$ defined as in (4.1) below. If there exist a complex structure $I$ and paracomplex structures $J_{1}$, $J_{2}$ of $E$ such that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$, then $r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ ($r\in\mbox{\boldmath{$R$}}\setminus\{0\}$, $\theta\in[0,2\pi)$) are $H$-nilpotent structures of $(E,h,\nabla)$ (Theorem 4.4). See [5], [14] for paraquaternionic structures, and see [10], [17] for neutral hyperKähler $4$-manifolds. An $H$-nilpotent structure $N$ of $(E,h,\nabla)$ defines a unique light-like subbundle $\pi^{*}_{N}$ of $(E,h)$ of rank $2n$ satisfying $E=\pi_{N}\oplus\pi^{*}_{N}$, and a unique $H$-nilpotent structure $N^{*}$ of $(E,h,\nabla)$ (the dual $H$-nilpotent structure of $N$) satisfying $\pi_{N^{*}}=\pi^{*}_{N}$ (Theorem 4.5). Based on this, we observe that if there exists an $H$-nilpotent structure $N$ of $(E,h,\nabla)$, then $h$, $\nabla$, $I:=(1/2)(N+N^{*})$, $J_{1}:=-IJ_{2}$, $J_{2}:=(1/2)(N-N^{*})$ form a neutral hyperKähler structure of $E$ (Corollary 4.6).
Remark 1.1
In [3], the author studied an $h$-reversing paracomplex structure $J$ of $(E,h)$ such that $\nabla J$ is locally represented as $\nabla J=\alpha\otimes N$ for a nowhere zero $1$-form $\alpha$ and a nilpotent structure $N$ related to $J$. The conformal Gauss maps of time-like minimal surfaces in $E^{3}_{1}$ give such paracomplex structures of the pull-back bundles. An oriented neutral $4n$-manifold with an almost paracomplex structure as above is considered to be a Walker manifold ([3]). We can find examples of almost paracomplex structures of $E^{4n}_{2n}$ as above ([3]). If $n=1$, then an $h$-reversing paracomplex structure of $(E,h)$ as above corresponds to a section of one of the time-like twistor spaces associated with $(E,h)$ such that the covariant derivative is fully light-like ([3]). We can obtain all the pairs of $h$-reversing almost paracomplex structures of $E^{4}_{2}$ such that each pair gives sections of the two time-like twistor spaces with fully light-like covariant derivatives ([3]).
This work was supported by JSPS KAKENHI Grant Number JP21K03228.
2 Nilpotent structures
Let $M$ be a manifold. Let $E$ be an oriented vector bundle over $M$ of rank $4n$ and $h$ a neutral metric of $E$. Let $N$ be a section of ${\rm End}\,E$. We call $N$ a nilpotent structure of $(E,h)$ if on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ satisfying $Ne=e\Lambda_{n}$ with
$\Lambda_{n}:=\left[\begin{array}[]{cccc}O_{n}&-I_{n}&O_{n}&I_{n}\\ I_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&I_{n}&O_{n}&-I_{n}\\ I_{n}&O_{n}&I_{n}&O_{n}\end{array}\right],$
where $I_{n}$ is the $n\times n$ unit matrix, $O_{n}$ is the $n\times n$ zero matrix, and we always suppose that $e_{1},\dots,e_{2n}$ are space-like and that $e_{2n+1},\dots,e_{4n}$ are time-like. Let $N$ be a nilpotent structure of $(E,h)$. For $\varepsilon=+$ or --, we call $N$ an $\varepsilon$-nilpotent structure of $(E,h)$ if on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e$ of $(E,h)$ giving the orientation of $E$ and satisfying $NeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\Lambda_{n}$, where
$I^{\prime}_{4n,\varepsilon}:=\left[\begin{array}[]{cccc}I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&I_{n}&O_{n}&O_{n}\\ O_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&I_{n,\varepsilon}\end{array}\right],$
and
$I_{1,\pm}:=\pm 1,\quad I_{n,\pm}:=\left[\begin{array}[]{cccc}\pm 1&0&\cdots&0\\ 0&1&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&1\end{array}\right]\ (n\geq 2).$
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $N$. Let $f=(f_{1},\dots,f_{2n},f_{2n+1},\dots,f_{4n})$ be an ordered pseudo-orthonormal local frame field of $(E,h)$ giving the orientation of $E$. Then $f$ is an admissible frame field of $N$ if and only if for each admissible frame field $e$ of $N$, an $SO(2n,2n)$-valued function $A$ on the intersection of the domains of $e$ and $f$ given by $fI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}A$ is valued in the Lie subgroup $G$ of $SO(2n,2n)$ defined by $A_{0}\Lambda_{n}=\Lambda_{n}A_{0}$ for $A_{0}\in SO(2n,2n)$ ([3]). For an admissible frame field $e$ of $N$, we set
$\begin{array}[]{lcl}\xi_{1}:=e_{1}-e_{2n+1},&&\xi_{i}:=e_{i}-e_{2n+i},\\ \xi_{n+1}:=e_{n+1}+\varepsilon e_{3n+1},&&\xi_{n+i}:=e_{n+i}+e_{3n+i}\end{array}\ (i=2,\dots,n)$ (2.1)
and $\xi=\xi_{1}\wedge\dots\wedge\xi_{2n}$. Then $\xi$ does not depend on the choice of an admissible frame field $e$ of $N$ ([3]). Therefore $N$ gives a section $\xi_{N}$ of the $2n$-fold exterior power $\bigwedge^{2n}\!E$ of $E$. A section $N$ of ${\rm End}\,E$ is a nilpotent structure of $(E,h)$ if and only if $N$ satisfies
  • (a)
    ${\rm Im}\,N={\rm Ker}\,N$, and $\pi_{N}:={\rm Im}\,N={\rm Ker}\,N$ is a subbundle of $E$ of rank $2n$ such that each fiber is light-like,
  • (b)
    $h(\phi,N\phi)=0$ for any local section $\phi$ of $E$
([2], [3]). In particular, $N$ gives a null structure on each fiber of $E$ and $h$ is null-Hermitian with respect to $N$ (see [11]). The section $\xi_{N}$ of $\bigwedge^{2n}\!E$ given by $N$ corresponds to the subbundle $\pi_{N}$ with the orientation assigned by $\xi_{N}$.
Remark 2.1
Suppose $n=1$. The $2$-fold exterior power $\bigwedge^{2}\!E$ of $E$ is a vector bundle over $M$ of rank $6$ and $h$ induces a metric $\hat{h}$ of $\bigwedge^{2}\!E$ of signature (2,4). In addition, $\bigwedge^{2}\!E$ is decomposed as $\bigwedge^{2}\!E=\bigwedge^{2}_{+}\!E\oplus\bigwedge^{2}_{-}\!E$ by two subbundles +$\bigwedge^{2}_{+}\!E$, $\bigwedge^{2}_{-}\!E$ of rank $3$ and the restriction of $\hat{h}$ on each of them has signature (1,2). The light-like twistor spaces associated with $(E,h)$ are fiber bundles $U_{0}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ in ±$\bigwedge^{2}_{\pm}\!E$ respectively such that each fiber is a light cone. Each light-like line subbundle of +$\bigwedge^{2}_{+}\!E$ or $\bigwedge^{2}_{-}\!E$ corresponds to a light-like subbundle of $(E,h)$ of rank $2$ and each $\varepsilon$-nilpotent structure $N$ of $(E,h)$ corresponds to a section $\Omega_{N}$ of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ given by $(1/\sqrt{2})\xi_{N}$ ([2], [3]).
Remark 2.2
Suppose $n=1$. The space-like twistor spaces $U_{+}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ associated with $(E,h)$ are fiber bundles in ±$\bigwedge^{2}_{\pm}\!E$ respectively such that each fiber is a hyperboloid of two sheets. A section of $U_{+}\!\left(\textstyle\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponds to a complex structure of $E$ preserving $h$. See [1], [5] for the space-like twistor spaces. The time-like twistor spaces $U_{-}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ associated with $(E,h)$ are fiber bundles in ±$\bigwedge^{2}_{\pm}\!E$ respectively such that each fiber is a hyperboloid of one sheet. A section of $U_{-}\!\left(\textstyle\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponds to a paracomplex structure of $E$ reversing $h$. See [1], [15], [16] for the time-like twistor spaces. See [7], [12], [13] for the twistor spaces in the case $h$ is a Riemannian (i.e., positive-definite) metric, which are the prototypes of $U_{+}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$, $U_{-}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ and $U_{0}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$.
3 The Walker condition
Let $\nabla$ be an $h$-connection of $E$, i.e., a connection of $E$ satisfying $\nabla h=0$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. We say that $N$ satisfies the Walker condition with respect to $\nabla$ if for any local section $\psi$ of $\pi_{N}$, $\nabla\psi$ is a $1$-form valued in $\pi_{N}$. See [6], [8], [9], [11], [19] for Walker manifolds.
Let $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ be an admissible frame field of $N$. Let $\omega=[\omega^{i}_{j}]$ be the connection form of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$. Then we have $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. We represent $\omega$ as
$\omega=\left[\begin{array}[]{cccc}D_{11}&D_{12}&D_{13}&D_{14}\\ D_{21}&D_{22}&D_{23}&D_{24}\\ D_{31}&D_{32}&D_{33}&D_{34}\\ D_{41}&D_{42}&D_{43}&D_{44}\end{array}\right],$
where $D_{kl}$ is the $(k,l)$-block of $\omega$, which is an $n\times n$ matrix such that the components are given by
$\omega^{i}_{j}\quad(i=(k-1)n+1,\dots,kn,\ j=(l-1)n+1,\dots,ln).$
Since $\omega$ is a $1$-form valued in the Lie algebra of $SO(2n,2n)$, we have
$\begin{split}&{}^{t}\!D_{ii}=-D_{ii}\ (i=1,2,3,4),\quad{}^{t}\!D_{21}=-D_{12},\quad{}^{t}\!D_{43}=-D_{34},\\ &{}^{t}\!D_{31}=D_{13},\quad{}^{t}\!D_{41}=D_{14},\quad{}^{t}\!D_{32}=D_{23},\quad{}^{t}\!D_{42}=D_{24}.\end{split}$ (3.1)
In addition, $N$ satisfies the Walker condition with respect to $\nabla$ if and only if on a neighborhood of each point of $M$, an admissible frame field $e$ of $N$ satisfies
$\begin{split}D_{11}-D_{13}+D_{31}-D_{33}&=O_{n},\\ D_{21}-D_{23}-D_{41}+D_{43}&=O_{n},\\ D_{22}+D_{24}-D_{42}-D_{44}&=O_{n}.\end{split}$ (3.2)
We denote by $\hat{\nabla}$ the connection of $\bigwedge^{2n}\!E$ induced by $\nabla$. We will prove
Proposition 3.1
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then the following hold::::
  • (a)
    $N$ satisfies the Walker condition with respect to $\nabla$ if and only if ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ for a $1$-form $\alpha$ on $M;$
  • (b)
    if $N$ is parallel with respect to $\nabla$, then ^$\hat{\nabla}\xi_{N}=0$, and therefore $N$ satisfies the Walker condition with respect to $\nabla$.
Proof  If an $\varepsilon$-nilpotent structure $N$ of $(E,h)$ satisfies the Walker condition with respect to $\nabla$, then we have ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$, where $\alpha$ is a $1$-form on $M$ which is locally represented as
$\alpha:=\left(-\sum^{n}_{i=1}\omega^{2n+i}_{i}+\sum^{n}_{i=1}\omega^{3n+i}_{n+i}\right).$ (3.3)
If $N$ satisfies ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ for a $1$-form $\alpha$ on $M$, then each $\nabla\xi_{i}$ is a $1$-form valued in $\pi_{N}$, which means that $N$ satisfies the Walker condition. Hence we obtain (a) in Proposition 3.1. Suppose that $N$ satisfies $\nabla N=0$. Then we have $\omega\Lambda_{n}=\Lambda_{n}\omega$. This is rewritten into
$\begin{split}&D_{11}-D_{13}=D_{22}-D_{42}=D_{33}-D_{31}=D_{44}-D_{24},\\ &D_{12}+D_{14}=-D_{21}+D_{41}=-D_{23}+D_{43}=-D_{32}-D_{34}.\end{split}$ (3.4)
Therefore we obtain
$\begin{split}&(\nabla\xi_{1}\ \dots\ \nabla\xi_{n})\\ &=(\xi_{1}\ \dots\ \xi_{n})(D_{11}-D_{13})+(\xi_{n+1}\ \dots\ \xi_{2n})(D_{21}-D_{23}),\\ &(\nabla\xi_{n+1}\ \dots\ \nabla\xi_{2n})\\ &=(\xi_{1}\ \dots\ \xi_{n})(D_{12}+D_{14})+(\xi_{n+1}\ \dots\ \xi_{2n})(D_{22}+D_{24}).\end{split}$ (3.5)
Since
$D_{22}+D_{24}=-{}^{t}(D_{22}-D_{42})=-{}^{t}(D_{11}-D_{13})$
by (3.1) and (3.4), the sum of the $(i,i)$-components of $D_{22}+D_{24}$ and $D_{11}-D_{13}$ is zero for $i=1,\dots,n$. Therefore ^$\hat{\nabla}\xi_{N}=0$. Hence we obtain (b) in Proposition 3.1. $\square$
Remark 3.1
It is already known that $\nabla N=0$ means that $N$ satisfies the Walker condition with respect to $\nabla$ (see [11]).
Remark 3.2
Suppose $n=1$. Then $\hat{\nabla}$ gives connections of ±$\bigwedge^{2}_{\pm}\!E$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then the corresponding section of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ is represented as $\Omega_{N}=(1/\sqrt{2})\xi_{N}$, and $\nabla N=0$ is equivalent to ^$\hat{\nabla}\xi_{N}=0$ ([2]). Since (3.1) means
$\omega^{i}_{i}=0\ (i=1,2,3,4),\quad\omega^{i+1}_{i}=-\omega^{i}_{i+1}\ (i=1,3),\quad\omega^{i}_{j}=\omega^{j}_{i}\ (|i-j|>1),$ (3.6)
the Walker condition is given by $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$ for an admissible frame field $e$ of $N$. In addition, noticing (3.3), we see that $N$ satisfies $\nabla N=0$ if and only if not only $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$ but also $\omega^{3}_{1}=\omega^{4}_{2}$ hold.
Remark 3.3
Suppose $n=1$. For a complex structure $I$ of $E$ preserving $h$, $\nabla I=0$ is equivalent to ^$\hat{\nabla}\Omega_{I}=0$ for the section $\Omega_{I}$ of $U_{+}\!\left(\textstyle\bigwedge^{2}_{+}\!E\right)$ or $U_{+}\!\left(\textstyle\bigwedge^{2}_{-}\!E\right)$ corresponding to $I$ ([1]). Similarly, for a paracomplex structure $J$ of $E$ reversing $h$, $\nabla J=0$ is equivalent to ^$\hat{\nabla}\Omega_{J}=0$ for the section $\Omega_{J}$ of $U_{-}\!\left(\textstyle\bigwedge^{2}_{+}\!E\right)$ or $U_{-}\!\left(\textstyle\bigwedge^{2}_{-}\!E\right)$ corresponding to $J$ ([1]).
Let $N$, $N^{\prime}$ be $\varepsilon$-nilpotent structures of $(E,h)$. Suppose that on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying
  • (i)
    $e$ is an admissible frame field of $N$,
  • (ii)
    $e^{\prime}:=(e_{1},\dots,e_{2n},-e_{2n+1},\dots,-e_{4n})$ is an admissible frame field of $N^{\prime}$.
Then $E=\pi_{N}\oplus\pi_{N^{\prime}}$. Referring to (3.2), we obtain
Proposition 3.2
Both of $N$ and $N^{\prime}$ satisfy the Walker condition if and only if the connection form of $\nabla$ with respect to $e$ as above satisfies not only (3.1) but also
$\begin{split}&D_{11}=D_{33},\quad D_{22}=D_{44},\quad D_{13}=D_{31},\quad D_{24}=D_{42},\\ &D_{43}=-D_{21},\quad D_{41}=-D_{23}.\end{split}$ (3.7)
Example 3.1
[…]
$F_{k}:=\left[\begin{array}[]{ccccc}a_{k}&b_{k}&0&\cdots&0\\ b_{k}&a_{k}&b_{k}&\ddots&\vdots\\ 0&b_{k}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&b_{k}\\ 0&\cdots&0&b_{k}&a_{k}\end{array}\right].$
Then we have $dF_{k}\wedge dF_{k}=O_{n}$ for $k=1,2$. We set
$\omega=\left[\begin{array}[]{cccc}O_{n}&O_{n}&dF_{1}&O_{n}\\ O_{n}&O_{n}&O_{n}&dF_{2}\\ dF_{1}&O_{n}&O_{n}&O_{n}\\ O_{n}&dF_{2}&O_{n}&O_{n}\end{array}\right].$
Then we have $d\omega+\omega\wedge\omega=O_{4n}$. Since we suppose that $\nabla$ is flat, there exists an ordered pseudo-orthonormal frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. Then $\omega$ is the connection form of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ and it satisfies (3.1) and (3.7). Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$, and $N^{\prime}$ an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e^{\prime}$ is an admissible frame field of $N^{\prime}$. Then $N$, $N^{\prime}$ satisfy the Walker condition. We have
^$\hat{\nabla}\xi_{N}=n(-da_{1}+da_{2})\otimes\xi_{N},\quad\hat{\nabla}\xi_{N^{\prime}}=n(da_{1}-da_{2})\otimes\xi_{N^{\prime}}.$
Therefore if $da_{1}\not=da_{2}$, then ^$\hat{\nabla}\xi_{N}$, ^$\hat{\nabla}\xi_{N^{\prime}}$ do not become zero and therefore by (b) in Proposition 3.1, $N$, $N^{\prime}$ are not parallel with respect to $\nabla$.
Example 3.2
Let $M$, $\nabla$ be as in Example 3.1. Let $a_{1}$, $a_{2}$ be functions on $M$ satisfying $da_{1}\not=da_{2}$. Suppose $n=1$ and set
$\omega=\left[\begin{array}[]{cccc}0&0&da_{1}&0\\ 0&0&0&da_{2}\\ da_{1}&0&0&0\\ 0&da_{2}&0&0\end{array}\right].$
Then we have $d\omega+\omega\wedge\omega=O_{4}$. Therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},e_{2},e_{3},e_{4})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4,\varepsilon}=eI^{\prime}_{4,\varepsilon}\omega$. Then $\omega$ is the connection form of $\nabla$ with respect to $eI^{\prime}_{4,\varepsilon}$ and it satisfies (3.1) and (3.7). Let $N$, $N^{\prime}$ be as in Example 3.1. Then $N$, $N^{\prime}$ satisfy the Walker condition. We have
^$\hat{\nabla}\xi_{N}=(-da_{1}+da_{2})\otimes\xi_{N},\quad\hat{\nabla}\xi_{N^{\prime}}=(da_{1}-da_{2})\otimes\xi_{N^{\prime}}.$
Therefore ^$\hat{\nabla}\xi_{N}$, ^$\hat{\nabla}\xi_{N^{\prime}}$ do not become zero and therefore $N$, $N^{\prime}$ are not parallel with respect to $\nabla$.
Example 3.3
Let $M$, $\nabla$ be as in Example 3.1 and suppose $n=1$. Let $N$ be an $\varepsilon$-nilpotent structure of $E$ with the Walker condition and $e$ an admissible frame field of $N$ defined on $M$. Then the connection form $\omega=[\omega^{i}_{j}]$ of $\nabla$ with respect to $eI^{\prime}_{4,\varepsilon}$ satisfies $d\omega+\omega\wedge\omega=0$ and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$. We can rewrite $d\omega+\omega\wedge\omega=0$ into
$\begin{split}d\omega^{2}_{1}+\omega^{3}_{2}\wedge\omega^{3}_{1}+\omega^{4}_{2}\wedge\omega^{4}_{1}&=0,\\ d\omega^{3}_{1}+\omega^{3}_{2}\wedge\omega^{2}_{1}-\omega^{4}_{3}\wedge\omega^{4}_{1}&=0,\\ d\omega^{4}_{1}+\omega^{4}_{2}\wedge\omega^{2}_{1}+\omega^{4}_{3}\wedge\omega^{3}_{1}&=0,\\ d\omega^{3}_{2}-\omega^{3}_{1}\wedge\omega^{2}_{1}-\omega^{4}_{3}\wedge\omega^{4}_{2}&=0,\\ d\omega^{4}_{2}-\omega^{4}_{1}\wedge\omega^{2}_{1}+\omega^{4}_{3}\wedge\omega^{3}_{2}&=0,\\ d\omega^{4}_{3}+\omega^{4}_{1}\wedge\omega^{3}_{1}+\omega^{4}_{2}\wedge\omega^{3}_{2}&=0.\end{split}$ (3.8)
Using the second and fifth equations in (3.8) and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$, we find a function $a$ satisfying
$da=\omega^{3}_{1}-\omega^{4}_{2};$ (3.9)
using the first and sixth equations in (3.8) and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$, we obtain
$d(\omega^{2}_{1}+\omega^{4}_{3})+(\omega^{2}_{1}+\omega^{4}_{3})\wedge da=0$ (3.10)
(using the third and fourth equations in (3.8) and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$, we can obtain (3.10) again). From (3.10), we see that there exists a function $b$ satisfying
$\omega^{2}_{1}+\omega^{4}_{3}=\omega^{4}_{1}+\omega^{3}_{2}=e^{a}db.$ (3.11)
From (3.9) and (3.11), we obtain
$\omega^{3}_{2}=e^{a}db-\omega^{4}_{1},\quad\omega^{4}_{3}=e^{a}db-\omega^{2}_{1},\quad\omega^{4}_{2}=\omega^{3}_{1}-da.$ (3.12)
Applying (3.12) to (3.8), we obtain
$\begin{split}&d\omega^{2}_{1}+2\omega^{3}_{1}\wedge\omega^{4}_{1}+e^{a}db\wedge\omega^{3}_{1}-da\wedge\omega^{4}_{1}=0,\\ &d\omega^{3}_{1}+2\omega^{2}_{1}\wedge\omega^{4}_{1}+e^{a}db\wedge(\omega^{2}_{1}-\omega^{4}_{1})=0,\\ &d\omega^{4}_{1}-2\omega^{2}_{1}\wedge\omega^{3}_{1}-da\wedge\omega^{2}_{1}+e^{a}db\wedge\omega^{3}_{1}=0.\end{split}$ (3.13)
Therefore the connection form $\omega=[\omega^{i}_{j}]$ satisfies (3.13) for functions $a$, $b$ on $M$. Conversely, if $1$-forms $\omega^{2}_{1}$, $\omega^{3}_{1}$, $\omega^{4}_{1}$ and functions $a$, $b$ on $M$ satisfy (3.13), then $\omega=[\omega^{i}_{j}]$ defined by (3.6) and (3.12) satisfies $d\omega+\omega\wedge\omega=0$ and therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},e_{2},e_{3},e_{4})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4,\varepsilon}=eI^{\prime}_{4,\varepsilon}\omega$. Then $e$ defines an $\varepsilon$-nilpotent structure $N$ of $E$ such that $e$ is an admissible frame field of $N$. Since $\omega^{i}_{j}$, $a$, $b$ satisfy (3.11), $N$ satisfies the Walker condition. In particular,
  • (a)
    $a$ is constant if and only if $\nabla N=0$, that is, ^$\hat{\nabla}\xi_{N}=0$;
  • (b)
    $b$ is constant if and only if there exists an $\varepsilon$-nilpotent structure $N^{\prime}$ of $E$ satisfying the Walker condition such that $e^{\prime}$ is an admissible frame field of $N^{\prime}$.
Suppose that $b$ is constant. Then (3.13) is represented as
$\begin{split}&d\omega^{2}_{1}+2\omega^{3}_{1}\wedge\omega^{4}_{1}-da\wedge\omega^{4}_{1}=0,\\ &d\omega^{3}_{1}+2\omega^{2}_{1}\wedge\omega^{4}_{1}=0,\\ &d\omega^{4}_{1}-2\omega^{2}_{1}\wedge\omega^{3}_{1}-da\wedge\omega^{2}_{1}=0.\end{split}$ (3.14)
If $\omega^{3}_{1}=dc$ for a function $c$ on $M$, then
$\omega^{2}_{1}=\dfrac{1}{2}(e^{a-2c}d\phi+e^{-a+2c}d\psi),\quad\omega^{4}_{1}=\dfrac{1}{2}(e^{a-2c}d\phi-e^{-a+2c}d\psi)$ (3.15)
for functions $\phi$, $\psi$ on $M$ satisfying $d\phi\wedge d\psi=0$. Conversely, for functions $c$, $\phi$, $\psi$ on $M$ with $d\phi\wedge d\psi=0$, $1$-forms $\omega^{2}_{1}$, $\omega^{4}_{1}$ defined as in (3.15) and $\omega^{3}_{1}:=dc$ satisfy (3.14) (see Example 3.2 for the case where $\phi=\psi=0$ and $da\not=0$).
Example 3.4
Suppose $n\geq 2$. Let $M$, $\nabla$ be as in Example 3.1. Let $a_{1},\dots,a_{n},b_{1},\dots,b_{n}$ be functions on $M$ satisfying
$\sum^{n}_{i=1}a_{i}=\sum^{n}_{i=1}b_{i}.$ (3.16)
We set
$A:=\left[\begin{array}[]{ccccc}a_{1}&0&\cdots&0\\ 0&a_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&a_{n}\end{array}\right],\quad B:=\left[\begin{array}[]{ccccc}b_{1}&0&\cdots&0\\ 0&b_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&b_{n}\end{array}\right]$
and
$\omega=\left[\begin{array}[]{cccc}O_{n}&O_{n}&dA&O_{n}\\ O_{n}&O_{n}&O_{n}&dB\\ dA&O_{n}&O_{n}&O_{n}\\ O_{n}&dB&O_{n}&O_{n}\end{array}\right].$
Then we have $d\omega+\omega\wedge\omega=O_{4n}$. Therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$, and $N^{\prime}$ an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e^{\prime}$ is an admissible frame field of $N^{\prime}$. Then $N$, $N^{\prime}$ satisfy the Walker condition, by (3.7). By (3.3) and (3.16), we have ^$\hat{\nabla}\xi_{N}=0$ and ^$\hat{\nabla}\xi_{N^{\prime}}=0$. If $dA\not=dB$, then by (3.4), $N$, $N^{\prime}$ are not parallel with respect to $\nabla$.
Example 3.5
Let $M$, $\nabla$ be as in Example 3.1. Let $F$ be an $n\times n$ symmetric matrix such that each component is a function on $M$ and suppose $dF\not=O_{n}$ and $dF\wedge dF=O_{n}$. We set
$\omega=\left[\begin{array}[]{cccc}O_{n}&O_{n}&O_{n}&dF\\ O_{n}&O_{n}&dF&O_{n}\\ O_{n}&dF&O_{n}&O_{n}\\ dF&O_{n}&O_{n}&O_{n}\end{array}\right].$ (3.17)
Then we have $d\omega+\omega\wedge\omega=O_{4n}$. Therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$. Then $N$ does not satisfy the Walker condition, because $\omega$ in (3.17) does not satisfy (3.2). We have
$(\nabla N)eI^{\prime}_{4n,\varepsilon}=2eI^{\prime}_{4n,\varepsilon}\left[\begin{array}[]{cccc}O_{n}&O_{n}&dF&O_{n}\\ O_{n}&O_{n}&O_{n}&-dF\\ dF&O_{n}&O_{n}&O_{n}\\ O_{n}&-dF&O_{n}&O_{n}\end{array}\right].$
In particular, if $dF=(1/2)dfI_{n}$ for a function $f$, then $\nabla N=df\otimes J$, where $J$ is a paracomplex structure of $E$ defined by
$JeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\left[\begin{array}[]{cccc}O_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\\ I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&-I_{n}&O_{n}&O_{n}\end{array}\right].$
In addition, if we suppose $n=1$, then $\nabla N=df\otimes J$ corresponds to the relation ^$\hat{\nabla}\Omega_{N}=df\otimes\Omega_{J}$, where $\Omega_{N}$ is a section of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $N$ and given by $(1/\sqrt{2})\xi_{N}$, and $\Omega_{J}$ is a section of $U_{-}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $J$.
4 Special nilpotent structures
Let $E$ be an oriented vector bundle over $M$ of rank $4n$ and $h$ a neutral metric of $E$.
Definition 4.1
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then for a Lie subgroup $K$ of $SO(2n,2n)$, $N$ is called a $K$-nilpotent structure of $(E,h)$ if there exist an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of admissible frame fields of $N$ satisfying
  • (i)
    each $e_{\lambda}$ is defined on $U_{\lambda}$,
  • (ii)
    if $U_{\lambda}\cap U_{\mu}\not=\emptyset$ for $\lambda$, $\mu\in\Lambda$, then $e_{\mu}I^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$ for a function $A_{\lambda\mu}$ valued in $K$.
Let $N$ be a $K$-nilpotent structure of $(E,h)$. By definition, $N$ gives a principal $K$-bundle $P$ over $M$ associated with $E$ such that each $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is a local section of $P$ (see [18] for principal fiber bundles). Let $e$ be an admissible frame field of $N$ defined on an open set $U$ of $M$. Then $e$ is called a $K$-admissible frame field if $eI^{\prime}_{4n,\varepsilon}$ gives a local section of $P$, that is, if for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, there exists a $K$-valued function $A_{\lambda}$ on $U\cap U_{\lambda}$ satisfying $eI^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda}$. Based on Definition 4.1, we obtain
Proposition 4.1
Any $\varepsilon$-nilpotent structure of $(E,h)$ is a $G$-nilpotent structure of $(E,h)$.
Let $\nabla$ be an $h$-connection of $E$.
Definition 4.2
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then for a Lie subgroup $K$ of $SO(2n,2n)$, $N$ is called a $K$-nilpotent structure of $(E,h,\nabla)$ if there exist an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 and
  • (iii)
    the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $K$.
Let $N$ be a $K$-nilpotent structure of $(E,h,\nabla)$. Then $\nabla$ gives a connection in the principal $K$-bundle $P$. Let $e$ be a $K$-admissible frame field of $N$. Then the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ is given by $\omega=A^{-1}_{\lambda}\omega_{\lambda}A_{\lambda}+A^{-1}_{\lambda}dA_{\lambda}$. Therefore $\omega$ is valued in the Lie algebra of $K$.
Suppose that $\nabla$ is flat, that is, the curvature tensor of $\nabla$ vanishes. Then there exists an ordered pseudo-orthonormal local frame field $e$ of $(E,h)$ on a neighborhood $U$ of each point of $M$ giving the orientation of $E$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$. Then the connection form of $\nabla$ with respect to $e$ vanishes. Therefore for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, if we define an $SO(2n,2n)$-valued function $\tilde{A}_{\lambda}$ on $U\cap U_{\lambda}$ by $eI^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}\tilde{A}_{\lambda}$, then $\tilde{A}_{\lambda}$ satisfies $d\tilde{A}_{\lambda}=-\omega_{\lambda}\tilde{A}_{\lambda}$. This relation means that $d\tilde{A}_{\lambda}$ is valued in the distribution $\mathscr{D}$ on $SO(2n,2n)$ obtained by the right translations of the Lie algebra of $K$. If we suppose that $U\cap U_{\lambda}$ is connected, then there exists an integral manifold of $\mathscr{D}$ containing $\tilde{A}_{\lambda}$. We can choose $e$ so that $\tilde{A}_{\lambda}$ is valued in $K$ at a point. Since $K$ is an integral manifold of $\mathscr{D}$, $\tilde{A}_{\lambda}$ is valued in $K$ over its domain. If we suppose $K\subset G$, then $e$ is a $K$-admissible frame field of $N$. Hence we obtain
Proposition 4.2
Suppose that $\nabla$ is flat. Let $K$ be a Lie subgroup of $G$. Then for a $K$-nilpotent structure of $(E,h,\nabla)$, there exists a $K$-admissible frame field $e$ of $N$ on a neighborhood of each point of $M$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$.
Remark 4.1
Let $\eta_{1},\dots,\eta_{n},\eta_{n+1},\dots,\eta_{2n}$ be $2n$ vectors of $E^{4n}_{2n}$ defined by
$\Lambda_{n}=[\eta_{n+1}\ \dots\ \eta_{2n}\ -\eta_{1}\ \dots\ -\eta_{n}\ \eta_{n+1}\ \dots\ \eta_{2n}\ \eta_{1}\ \dots\ \eta_{n}].$
Let $W$ be a $2n$-dimensional subspace of $E^{4n}_{2n}$ spanned by $\eta_{1},\dots,\eta_{n},\eta_{n+1},\dots,\eta_{2n}$. Then $W$ is a light-like subspace of $E^{4n}_{2n}$. Let $SO(2n,2n)_{W}$ be a Lie subgroup of $SO(2n,2n)$ which consists of elements preserving $W$. Then we have $G\subset SO(2n,2n)_{W}$ and $G\not=SO(2n,2n)_{W}$ ([3]). A $SO(2n,2n)_{W}$-nilpotent structure of $(E,h,\nabla)$ is just an $\varepsilon$-nilpotent structure of $(E,h)$ satisfying the Walker condition. If $K=SO(2n,2n)_{W}$, then the conclusion of Proposition 4.2 is not necessarily valid, because of Example 3.1 $\sim$ Example 3.4, while for an almost $\varepsilon$-nilpotent structure $N$ of $E^{4n}_{2n}$ satisfying the Walker condition with respect to the Levi-Civita connection, $\pi_{N}$ gives a constant $2n$-dimensional light-like subspace of $E^{4n}_{2n}$.
We will prove
Proposition 4.3
A $G$-nilpotent structure of $(E,h,\nabla)$ is just an $\varepsilon$-nilpotent structure of $(E,h)$ parallel with respect to $\nabla$.
Proof  Let $N$ be a $G$-nilpotent structure of $(E,h,\nabla)$. Then the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ for any admissible frame field $e$ of $N$ is valued in the Lie algebra of $G$. Therefore we have $\omega\Lambda_{n}=\Lambda_{n}\omega$. By this, we obtain $\nabla(Ne)=N\nabla e$. Therefore $N$ is parallel with respect to $\nabla$. Suppose that $N$ is an $\varepsilon$-nilpotent structure of $(E,h)$ parallel with respect to $\nabla$. Then we have $\omega\Lambda_{n}=\Lambda_{n}\omega$ for the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ for any admissible frame field $e$ of $N$. This means that $\omega$ is valued in the Lie algebra of $G$. Therefore we can find an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 and (iii) in Definition 4.2 for $K=G$. Therefore $N$ is a $G$-nilpotent structure of $(E,h,\nabla)$. $\square$
Let $H$ be a Lie subgroup of $SO(2n,2n)$ defined by
$H:=\left\{\left[\begin{array}[]{cccc}A_{11}&-A_{21}&A_{31}&A_{41}\\ A_{21}&A_{11}&-A_{41}&A_{31}\\ A_{31}&-A_{41}&A_{11}&A_{21}\\ A_{41}&A_{31}&-A_{21}&A_{11}\end{array}\right]\in SO(2n,2n)\right\}.$ (4.1)
Then $H$ is a Lie subgroup of $G$. If $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then for an $H$-admissible frame field $e$ of $N$, the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ is represented as
$\omega=\left[\begin{array}[]{cccc}D_{11}&-D_{21}&D_{31}&D_{41}\\ D_{21}&D_{11}&-D_{41}&D_{31}\\ D_{31}&-D_{41}&D_{11}&D_{21}\\ D_{41}&D_{31}&-D_{21}&D_{11}\end{array}\right]\qquad\left(\begin{array}[]{l}{}^{t}D_{11}=-D_{11},\\ {}^{t}D_{21}=D_{21},\\ {}^{t}D_{31}=D_{31},\\ {}^{t}D_{41}=D_{41}\end{array}\right).$ (4.2)
The main objects of study in this section are $H$-nilpotent structures of $(E,h,\nabla)$.
Let $I$ be a complex structure of $E$ and $J_{1}$, $J_{2}$ paracomplex structures of $E$. Suppose that $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$, that is, suppose that $I$, $J_{1}$, $J_{2}$ satisfy
  • (i)
    $IJ_{1}=J_{2}$
and span a subbundle $E_{I,J_{1},J_{2}}$ of ${\rm End}\,E$ of rank $3$. In addition, suppose that $h$ is adapted to $E_{I,J_{1},J_{2}}$, that is, suppose
  • (ii)
    $I$ preserves $h$ and $J_{1}$, $J_{2}$ reverse $h$, that is, $h$ is Hermitian with respect to $I$ and paraHermitian with respect to $J_{1}$, $J_{2}$.
See [5], [14] for paraquaternionic structures (notice the signature in (i)). We say that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$ if $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ satisfy (i), (ii) and
  • (iii)
    $I$, $J_{1}$, $J_{2}$ are parallel with respect to $\nabla$.
See [10], [17] for neutral hyperKähler $4$-manifolds.
We will prove
Theorem 4.4
The following hold::::
  • (a)
    a section of $E_{I,J_{1},J_{2}}$ given by $N_{r,\theta}:=r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ for $r\in\mbox{\boldmath{$R$}}\setminus\{0\}$ and $\theta\in[0,2\pi)$ is an $H$-nilpotent structure of $(E,h);$
  • (b)
    if $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$, then $N_{r,\theta}$ is an $H$-nilpotent structure of $(E,h,\nabla)$.
Proof  Let $\{U_{\lambda}\}_{\lambda\in\Lambda}$ be an open cover of $M$ such that on each $U_{\lambda}$, there exists an ordered pseudo-orthonormal local frame field $e_{\lambda}=(e_{\lambda,1},\dots,e_{\lambda,2n},e_{\lambda,2n+1},\dots,e_{\lambda,4n})$ of $E$ satisfying
$Ie_{\lambda,i}=e_{\lambda,n+i},\quad J_{1}e_{\lambda,i}=e_{\lambda,2n+i},\quad J_{2}e_{\lambda,i}=e_{\lambda,3n+i}\quad(i=1,\dots,n).$ (4.3)
If $U_{\lambda}\cap U_{\mu}\not=\emptyset$ for $\lambda$, $\mu\in\Lambda$ and if we represent $e_{\mu}$ as $e_{\mu}=e_{\lambda}A_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$, then $A_{\lambda\mu}$ is valued in $H$. Hence $I$, $J_{1}$, $J_{2}$ define a principal $H$-bundle $P$ over $M$ associated with $E$ so that each $e_{\lambda}$ is a local section of $P$. For each $\lambda\in\Lambda$, let $N_{\lambda}$ be a section of the restriction of ${\rm End}\,E$ on $U_{\lambda}$ defined by $N_{\lambda}e_{\lambda}=e_{\lambda}\Lambda_{n}$. Then noticing that $H$ is contained in $G$, we can define a section $N$ of ${\rm End}\,E$ by $N=N_{\lambda}$ on each $U_{\lambda}$, which is an $\varepsilon$-nilpotent structure of $(E,h)$ for $\varepsilon=+$ or -- according to whether $e_{\lambda}$ gives the orientation of $E$ or not. In addition, $N$ is an $H$-nilpotent structure of $(E,h)$ and for each $\lambda\in\Lambda$, $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is an $H$-admissible frame field of $N$. By (4.3), $N$ coincides with $I+J_{2}=N_{1,0}$. For $\theta\in[0,2\pi)$, we set
$J^{\prime}_{1}:=(\cos\theta)J_{1}+(\sin\theta)J_{2},\quad J^{\prime}_{2}:=-(\sin\theta)J_{1}+(\cos\theta)J_{2}.$
Then $J^{\prime}_{1}$, $J^{\prime}_{2}$ are paracomplex structures of $E$ reversing $h$, and $I$, $J^{\prime}_{1}$, $J^{\prime}_{2}$ satisfy $IJ^{\prime}_{1}=J^{\prime}_{2}$ and span $E_{I,J_{1},J_{2}}$. Therefore referring to the above discussion, we obtain an $H$-nilpotent structure of $(E,h)$, which is given by
$I+J^{\prime}_{2}=I-(\sin\theta)J_{1}+(\cos\theta)J_{2}=N_{1,\theta}.$
For $t\not=0$ and $\delta\in\{1,-1\}$, we set
$I^{\prime}:=\delta(({\rm cosh}\,t)I+({\rm sinh}\,t)J^{\prime}_{2}),\quad J^{\prime\prime}_{2}:=\delta(({\rm sinh}\,t)I+({\rm cosh}\,t)J^{\prime}_{2}).$
Then $I^{\prime}$ is a complex structure of $E$ preserving $h$ and $J^{\prime\prime}_{2}$ is a paracomplex structure of $E$ reversing $h$. In addition, $I^{\prime}$, $J^{\prime}_{1}$ and $J^{\prime\prime}_{2}$ satisfy $I^{\prime}J^{\prime}_{1}=J^{\prime\prime}_{2}$ and span $E_{I,J_{1},J_{2}}$. Therefore referring to the above discussion, we obtain an $H$-nilpotent structure of $(E,h)$ given by
$I^{\prime}+J^{\prime\prime}_{2}=\delta e^{t}(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})=N_{\delta e^{t},\theta}.$
Hence we obtain (a) in Theorem 4.4. We set
$\begin{split}(\nabla e_{\lambda,1}\ \dots\ \nabla e_{\lambda,n})=&(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{11}+(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{21}\\ &+(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{31}+(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{41}.\end{split}$
Suppose that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$. Then by $\nabla I=0$, we obtain
$\begin{split}(\nabla e_{\lambda,n+1}\ \dots\ \nabla e_{\lambda,2n})=&-(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{21}+(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{11}\\ &-(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{41}+(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{31}.\end{split}$
Similarly, by $\nabla J_{1}=0$, we obtain
$\begin{split}(\nabla e_{\lambda,2n+1}\ \dots\ \nabla e_{\lambda,3n})=&(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{31}-(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{41}\\ &+(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{11}-(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{21},\end{split}$
and by $\nabla J_{2}=0$, we obtain
$\begin{split}(\nabla e_{\lambda,3n+1}\ \dots\ \nabla e_{\lambda,4n})=&(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{41}+(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{31}\\ &+(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{21}+(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{11}.\end{split}$
Therefore the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}$ is given as in (4.2) and therefore the connection form $\omega_{\lambda}$ is valued in the Lie algebra of $H$. Hence $\nabla$ gives a connection in $P$ and $N=N_{1,0}$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Similarly, for $r\in\mbox{\boldmath{$R$}}\setminus\{0\}$ and $\theta\in[0,2\pi)$, $N_{r,\theta}$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Hence we obtain (b) in Theorem 4.4. $\square$
Remark 4.2
Whether $e_{\lambda}$ gives the orientation of $E$ does not depend on $\lambda\in\Lambda$, and $N_{r,\theta}$ is a $+$ or ---nilpotent structure according to whether $e_{\lambda}$ gives the orientation or not, as was already seen in the above proof.
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ and suppose that $N$ is an $H$-nilpotent structure of $(E,h)$. Then the light-like subbundle $\pi_{N}$ of rank $2n$ is determined by $N$ and locally spanned by $\xi_{1},\dots,\xi_{2n}$ as in (2.1) for an $H$-admissible frame field $e$ of $N$. In addition, we will prove
Theorem 4.5
Let $N$ be as above.
  • (a)
    There exists a unique light-like subbundle $\pi^{*}_{N}$ of $(E,h)$ of rank $2n$ which is locally spanned by
    $\begin{array}[]{lcl}\xi^{\prime}_{1}:=e_{1}+e_{2n+1},&&\xi^{\prime}_{i}:=e_{i}+e_{2n+i},\\ \xi^{\prime}_{n+1}:=e_{n+1}-\varepsilon e_{3n+1},&&\xi^{\prime}_{n+i}:=e_{n+i}-e_{3n+i}\end{array}\ (i=2,\dots,n)$ (4.4)
    for any $H$-admissible frame field $e$ of $N$. In particular, $E=\pi_{N}\oplus\pi^{*}_{N}$.
  • (b)
    There exists a unique $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{*}$ of $(E,h)$ such that each $H$-admissible frame field of $N^{*}$ is given by $e^{\prime}=(e_{1},\dots,e_{2n},-e_{2n+1},\dots,-e_{4n})$ for an $H$-admissible frame field $e$ of $N$. In particular, $\pi_{N^{*}}=\pi^{*}_{N}$.
  • (c)
    If $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then $N^{*}$ is also an $H$-nilpotent structure of $(E,h,\nabla)$.
Proof  For an $H$-admissible frame field $e$ of $N$, the subbundle of the restriction of $E$ on the domain of $e$ spanned by $\xi^{\prime}_{i}$, $\xi^{\prime}_{n+i}$ ($i=1,\dots,n$) as in (4.4) does not depend on the choice of $e$, because of the definition of $H$ in (4.1). Therefore there exists a unique light-like subbundle $\pi^{*}_{N}$ of $(E,h)$ of rank $2n$ satisfying (a) in Theorem 4.5. Hence we obtain (a) in Theorem 4.5. For $N$, we can find an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of $H$-admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 for $K=H$. We set
$A^{\prime}_{\lambda\mu}:=I_{2n,2n}A_{\lambda\mu}I_{2n,2n},\quad I_{2n,2n}:=\left[\begin{array}[]{cccc}I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&I_{n}&O_{n}&O_{n}\\ O_{n}&O_{n}&-I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\end{array}\right]$
for $\lambda$, $\mu\in\Lambda$ satisfying $U_{\lambda}\cap U_{\mu}\not=\emptyset$. Then $A^{\prime}_{\lambda\mu}$ satisfies $e^{\prime}_{\mu}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$, and since $A_{\lambda\mu}$ is valued in $H$, $A^{\prime}_{\lambda\mu}$ is also valued in $H$. Therefore there exists an $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{*}$ of $(E,h)$ such that $e^{\prime}_{\lambda}$ ($\lambda\in\Lambda$) are $H$-admissible frame fields of $N^{*}$ and then $N^{*}$ is uniquely determined by $\{e^{\prime}_{\lambda}\}_{\lambda\in\Lambda}$. Let $\tilde{e}$ be an $H$-admissible frame field of $N^{*}$ defined on an open set $U$ of $M$. Then for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, there exists an $H$-valued function $A^{\prime}_{\lambda}$ on $U\cap U_{\lambda}$ satisfying $\tilde{e}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda}$. This relation is equivalent to $\tilde{e}^{\prime}I^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda}$ with $A_{\lambda}:=I_{2n,2n}A^{\prime}_{\lambda}I_{2n,2n}$. Since $A_{\lambda}$ is $H$-valued, $\tilde{e}^{\prime}$ is an $H$-admissible frame field of $N$. Hence we obtain (b) in Theorem 4.5. Suppose that $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Then we can suppose that the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. Therefore the connection form $\omega^{\prime}_{\lambda}$ of $\nabla$ with respect to $e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. Hence we obtain (c) in Theorem 4.5. $\square$
Let $N$ be an $H$-nilpotent structure of $(E,h)$. Then we call $N^{*}$ as in Theorem 4.5 the dual $H$-nilpotent structure of $N$. We see that $N$ is the dual $H$-nilpotent structure of $N^{*}$: $(N^{*})^{*}=N$.
For an $H$-nilpotent structure $N$ of $(E,h)$, we set
$I:=\dfrac{1}{2}(N+N^{*}),\quad J_{2}:=\dfrac{1}{2}(N-N^{*}),\quad J_{1}:=-IJ_{2}.$ (4.5)
Then $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$. In addition, if $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$. Hence we obtain
Corollary 4.6
An $H$-nilpotent structure $N$ of $(E,h)$ defines a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$ by (4.5). In addition, if $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$.
Remark 4.3
Let $I$ be a complex structure of $E$ and $J_{1}$, $J_{2}$ paracomplex structures of $E$. Suppose that $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$. Then $N:=I+J_{2}$ is an $H$-nilpotent structure of $(E,h)$ and $N^{*}:=I-J_{2}$ is its dual $H$-nilpotent structure.
Remark 4.4
If $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then we see by (4.2) that the connection form of an $H$-admissible frame field $e$ of $N$ satisfies (3.7).
Remark 4.5
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Let $e$, $\tilde{e}$ be admissible frame fields of $N$ and $A$ a $G$-valued function on the intersection of the domains of $e$, $\tilde{e}$ given by $\tilde{e}I^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}A$. Then the subbundle spanned by $\xi_{i}$, $\xi_{n+i}$ ($i=1,\dots,n$) coincides with the subbundle spanned by $\tilde{\xi}_{i}$, $\tilde{\xi}_{n+i}$ ($i=1,\dots,n$). In addition, if the subbundle spanned by $\xi^{\prime}_{i}$, $\xi^{\prime}_{n+i}$ ($i=1,\dots,n$) coincides with the subbundle spanned by $\tilde{\xi}^{\prime}_{i}$, $\tilde{\xi}^{\prime}_{n+i}$ ($i=1,\dots,n$), then $A$ is valued in $H$.