### v3

# Nilpotent structures of oriented neutral vector bundles

Abstract  In this paper, we study nilpotent structures of an oriented vector bundle $E$ of rank $4n$ with a neutral metric $h$ and an $h$-connection $\nabla$. We define $H$-nilpotent structures of $(E,h,\nabla)$ for a Lie subgroup $H$ of $SO(2n,2n)$ related to neutral hyperKähler structures. We observe that there exist a complex structure $I$ and paracomplex structures $J_{1}$, $J_{2}$ of $E$ such that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$ if and only if there exists an $H$-nilpotent structure of $(E,h,\nabla)$.
1 Introduction
Nilpotent structures of an oriented neutral vector bundle $(E,h)$ of rank $4n$ are analogues of complex structures and paracomplex structures. In particular, if $n=1$, then a nilpotent structure corresponds to a section of one of the light-like twistor spaces associated with $(E,h)$ (see [3], [4], [5] for the light-like twistor spaces) and a complex (respectively, paracomplex) structure preserving (respectively, reversing) the neutral metric $h$ corresponds to a section of one of the space-like (respectively, time-like) twistor spaces associated with $(E,h)$ (see [2], [7] for the space-like twistor spaces and see [2], [22], [23] for the time-like twistor spaces). A nilpotent structure $N$ gives a null structure on each fiber of $E$ and $h$ is null-Hermitian with respect to $N$ (see [15]).
A nilpotent structure $N$ gives a light-like subbundle $\pi_{N}$ of $(E,h)$ of rank $2n$. In addition, $N$ gives a nowhere zero section $\xi_{N}$ of a line bundle $\bigwedge^{2n}\!\pi_{N}$, which is constructed by a local frame field of $\pi_{N}$ given by an admissible frame field of $N$. The fact that $\xi_{N}$ is well-defined is related to a Lie subgroup $G$ of $SO(2n,2n)$. The transition function between two admissible frame fields of $N$ is valued in $G$. It gives a local section of ${\rm End}\,\pi_{N}$ and its determinant is identically equal to one ([4]). In the present paper, we will see that $N$ gives a nondegenerate section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$ satisfying $\xi_{N}=((-1)^{\frac{n(n-1)}{2}}/n!)\Theta^{n}_{N}$ and that a nondegenerate section $\Theta$ of the $2$-fold exterior power of a light-like subbundle $L$ of $(E,h)$ of rank $2n$ gives a nilpotent structure $N$ of $(E,h)$ satisfying $\pi_{N}=L$ and $\Theta_{N}=\Theta$ (Theorem 2.1). If $n=1$, then the section of one of the light-like twistor spaces corresponding to $N$ is given by $(1/\sqrt{2})\xi_{N}$ ([3]).
Let $\nabla$ be an $h$-connection of $E$, i.e., a connection of $E$ satisfying $\nabla h=0$. Then the Walker condition of $N$ is defined, based on the definition of Walker manifolds (see [8], [11], [12], [15], [30] for Walker manifolds). The Walker condition of $N$ is characterized by ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ with a $1$-form $\alpha$ on $M$ for the connection $\hat{\nabla}$ of $\bigwedge^{2n}\!E$ induced by $\nabla$ ((a) in Proposition 3.1). Therefore the Walker condition just means that $\hat{\nabla}$ induces a connection of $\bigwedge^{2n}\!\pi_{N}$. If we also denote by $\hat{\nabla}$ the induced connection of $\bigwedge^{2}\!E$, then $\nabla N=0$ is equivalent to ^$\hat{\nabla}\Theta_{N}=0$, and therefore $\nabla N=0$ yields ^$\hat{\nabla}\xi_{N}=0$ ((b) in Proposition 3.1), which means that a nilpotent structure $N$ parallel with respect to $\nabla$ satisfies the Walker condition (see [15]). If $n=1$, then $\nabla N=0$ is equivalent to ^$\hat{\nabla}\xi_{N}=0$ ([3]). There exists a nilpotent structure $N$ satisfying the Walker condition and ^$\hat{\nabla}\xi_{N}\not=0$ (Example 3.1, Example 3.2, Example 3.3). In addition, if $n\geq 2$, then there exists a nilpotent structure $N$ satisfying ^$\hat{\nabla}\xi_{N}=0$ and $\nabla N\not=0$ (Example 3.4). Such an example $N$ in Example 3.1 $\sim$ Example 3.4 admits a nilpotent structure $N^{\prime}$ of $(E,h)$ satisfying the Walker condition and $E=\pi_{N}\oplus\pi_{N^{\prime}}$.
In the present paper, we define special nilpotent structures related to a Lie subgroup $K$ of $SO(2n,2n)$. Each of them gives a principal $K$-bundle $P$ associated with $E$ and special admissible frame fields we use are given by local sections of $P$. In addition, $\nabla$ gives a connection in $P$, so that the connection form of $\nabla$ with respect to such a local section is valued in the Lie algebra of $K$. We call such a special nilpotent structure a $K$-nilpotent structure of $(E,h,\nabla)$. If $\nabla$ is flat and if $K\subset G$, then for a $K$-nilpotent structure of $(E,h,\nabla)$, we can find a local section $e$ of $P$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$ (Proposition 4.2). A $G$-nilpotent structure of $(E,h,\nabla)$ is just a nilpotent structure of $(E,h)$ parallel with respect to $\nabla$ (Proposition 4.3). Let $H$ be a Lie subgroup of $G$ defined as in (4.1) below. If there exist a complex structure $I$ and paracomplex structures $J_{1}$, $J_{2}$ of $E$ such that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$, then $r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ ($r\in\mbox{\boldmath{$R$}}\setminus\{0\}$, $\theta\in[0,2\pi)$) are $H$-nilpotent structures of $(E,h,\nabla)$ (Theorem 4.4). See [7], [21] for paraquaternionic structures, and see [13], [25] for neutral hyperKähler $4$-manifolds. An $H$-nilpotent structure $N$ of $(E,h,\nabla)$ defines a unique light-like subbundle $\pi^{\times}_{N}$ of $(E,h)$ of rank $2n$ satisfying $E=\pi_{N}\oplus\pi^{\times}_{N}$, and a unique $H$-nilpotent structure $N^{\times}$ of $(E,h,\nabla)$ (the dual $H$-nilpotent structure of $N$) satisfying $\pi_{N^{\times}}=\pi^{\times}_{N}$ (Theorem 4.5). Based on this, we observe that if there exists an $H$-nilpotent structure $N$ of $(E,h,\nabla)$, then $h$, $\nabla$, $I:=(1/2)(N+N^{\times})$, $J_{1}:=-IJ_{2}$, $J_{2}:=(1/2)(N-N^{\times})$ form a neutral hyperKähler structure of $E$ (Corollary 4.6).
Remark 1.1
In [4], the author studied an $h$-reversing paracomplex structure $J$ of $(E,h)$ such that $\nabla J$ is locally represented as $\nabla J=\alpha\otimes N$ for a nowhere zero $1$-form $\alpha$ and a nilpotent structure $N$ related to $J$. The conformal Gauss maps of time-like minimal surfaces in $E^{3}_{1}$ give such paracomplex structures of the pull-back bundles. An oriented neutral $4n$-manifold with an almost paracomplex structure as above is considered to be a Walker manifold ([4]). We can find examples of almost paracomplex structures of $E^{4n}_{2n}$ as above ([4]). If $n=1$, then an $h$-reversing paracomplex structure of $(E,h)$ as above corresponds to a section of one of the time-like twistor spaces associated with $(E,h)$ such that the covariant derivative is fully light-like ([4]). We can obtain all the pairs of $h$-reversing almost paracomplex structures of $E^{4}_{2}$ such that each pair gives sections of the two time-like twistor spaces with fully light-like covariant derivatives ([4]).
Remark 1.2
The metrics of vector bundles in the present paper are neutral. We can refer to [18] for almost paracomplex structures on Riemannian or neutral 4-manifolds. See [14] for a characterization of anti-self-dual null-Kähler $4$-manifolds. Neutral metrics appear in the studies of spaces of oriented geodesics ([1], [20], [28], [29]), the ultra-hyperbolic equation ([6], [10], [19], [24]) and quantum field theories ([27]).
This work was supported by JSPS KAKENHI Grant Number JP21K03228.
2 Nilpotent structures
Let $M$ be a manifold. Let $E$ be an oriented vector bundle over $M$ of rank $4n$ and $h$ a neutral metric of $E$. Let $N$ be a section of ${\rm End}\,E$. We call $N$ a nilpotent structure of $(E,h)$ if on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ satisfying $Ne=e\Lambda_{n}$ with
$\Lambda_{n}:=\left[\begin{array}[]{cccc}O_{n}&-I_{n}&O_{n}&I_{n}\\ I_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&I_{n}&O_{n}&-I_{n}\\ I_{n}&O_{n}&I_{n}&O_{n}\end{array}\right],$
where $I_{n}$ is the $n\times n$ unit matrix, $O_{n}$ is the $n\times n$ zero matrix, and we always suppose that $e_{1},\dots,e_{2n}$ are space-like and that $e_{2n+1},\dots,e_{4n}$ are time-like. Let $N$ be a nilpotent structure of $(E,h)$. For $\varepsilon=+$ or --, we call $N$ an $\varepsilon$-nilpotent structure of $(E,h)$ if on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e$ of $(E,h)$ giving the orientation of $E$ and satisfying $NeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\Lambda_{n}$, where
$I^{\prime}_{4n,\varepsilon}:=\left[\begin{array}[]{cccc}I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&I_{n}&O_{n}&O_{n}\\ O_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&I_{n,\varepsilon}\end{array}\right],$
and
$I_{1,\pm}:=\pm 1,\quad I_{n,\pm}:=\left[\begin{array}[]{cccc}\pm 1&0&\cdots&0\\ 0&1&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&1\end{array}\right]\ (n\geq 2).$
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $N$. Let $f=(f_{1},\dots,f_{2n},f_{2n+1},\dots,f_{4n})$ be an ordered pseudo-orthonormal local frame field of $(E,h)$ giving the orientation of $E$. Then $f$ is an admissible frame field of $N$ if and only if for each admissible frame field $e$ of $N$, an $SO(2n,2n)$-valued function $A$ on the intersection of the domains of $e$ and $f$ given by $fI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}A$ is valued in the Lie subgroup $G$ of $SO(2n,2n)$ defined by $A_{0}\Lambda_{n}=\Lambda_{n}A_{0}$ for $A_{0}\in SO(2n,2n)$ ([4]). A section $N$ of ${\rm End}\,E$ is a nilpotent structure of $(E,h)$ if and only if $N$ satisfies
  • (a)
    ${\rm Im}\,N={\rm Ker}\,N$, and $\pi_{N}:={\rm Im}\,N={\rm Ker}\,N$ is a subbundle of $E$ of rank $2n$ such that each fiber is light-like,
  • (b)
    $h(\phi,N\phi)=0$ for any local section $\phi$ of $E$
([3], [4]). In particular, $N$ gives a null structure on each fiber of $E$ and $h$ is null-Hermitian with respect to $N$ (see [15]).
Let $E^{*}$ be the dual bundle of $E$. Then $h$ induces a neutral metric $h^{*}$ of $E^{*}$. For an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$, let $e^{*}:=(e^{1},\dots,e^{2n},e^{2n+1},\dots,e^{4n})$ be the dual frame field of $e$. Then $e^{*}$ is an ordered pseudo-orthonormal local frame field of $(E^{*},h^{*})$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then $N$ gives a section $N^{*}$ of ${\rm End}\,E^{*}$ by
$N^{*}\phi^{*}:=\phi^{*}\circ N$ (2.1)
for a local section $\phi^{*}$ of $E^{*}$. Let $e$ be an admissible frame field of $N$. Then $e^{*}$ satisfies $N^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n}$. By this, we obtain $h^{*}(\phi^{*},N^{*}\phi^{*})=0$ for any local section $\phi^{*}$ of $E^{*}$. Therefore, if we set
$\Theta_{N}(\phi^{*},\psi^{*}):=h^{*}(\phi^{*},N^{*}\psi^{*})$ (2.2)
for local sections $\phi^{*}$, $\psi^{*}$ of $E^{*}$, then $\Theta_{N}(\psi^{*},\phi^{*})=-\Theta_{N}(\phi^{*},\psi^{*})$. This means that $\Theta_{N}$ is a section of the $2$-fold exterior power $\bigwedge^{2}\!E$ of $E$ given by $N$. In addition, by $N^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n}$, we obtain
$\Theta_{N}=\sum^{n}_{i=1}\xi_{i}\wedge\xi_{n+i},$
where
$\begin{array}[]{lcl}\xi_{1}:=e_{1}-e_{2n+1},&&\xi_{i}:=e_{i}-e_{2n+i},\\ \xi_{n+1}:=e_{n+1}+\varepsilon e_{3n+1},&&\xi_{n+i}:=e_{n+i}+e_{3n+i}\end{array}\ (i=2,\dots,n).$ (2.3)
Therefore $\Theta_{N}$ is a section of $\bigwedge^{2}\!\pi_{N}$. In addition, $\Theta_{N}$ is nondegenerate, that is, $\Theta_{N}$ gives a nowhere zero section $\xi_{N}$ of $\bigwedge^{2n}\!\pi_{N}$ by
$\xi_{N}=\dfrac{(-1)^{\frac{n(n-1)}{2}}}{n!}\Theta^{n}_{N}=\xi_{1}\wedge\dots\wedge\xi_{2n}.$ (2.4)
Let $L$ be a subbundle of $E$ of rank $2n$. We say that $L$ is an $\varepsilon$-light-like subbundle of $(E,h)$ if on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ such that $\xi_{1},\dots,\xi_{n},\xi_{n+1},\dots,\xi_{2n}$ as in (2.3) form a local frame field of $L$. Let $L$ be an $\varepsilon$-light-like subbundle of $(E,h)$. Let $\Theta$ be a nondegenerate section of $\bigwedge^{2}\!L$. Then on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e$ of $(E,h)$ giving the orientation of $E$ and satisfying $\Theta=\sum^{n}_{i=1}\xi_{i}\wedge\xi_{n+i}$ with (2.3). For $\Theta$, let $N^{*}$ be a section of ${\rm End}\,E^{*}$ given by $\Theta(\phi^{*},\psi^{*})=h^{*}(\phi^{*},N^{*}\psi^{*})$ for local sections $\phi^{*}$, $\psi^{*}$ of $E^{*}$. Then $N^{*}$ satisfies $h^{*}(\phi^{*},N^{*}\phi^{*})=0$ and $N^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n}$. For $N^{*}$, let $N$ be a section of ${\rm End}\,E$ given by $N^{*}\phi^{*}=\phi^{*}\circ N$. Then $N$ is an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$ and $N$ satisfies $\pi_{N}=L$ and $\Theta_{N}=\Theta$. Hence we obtain
Theorem 2.1
There exists a one-to-one correspondence between the set of $\varepsilon$-nilpotent structures of $(E,h)$ and the set of nondegenerate sections of the $2$-fold exterior powers of $\varepsilon$-light-like subbundles of $(E,h)$:::: each $\varepsilon$-nilpotent structure $N$ corresponds to a nondegenerate section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$ by (2.1) and (2.2).
Remark 2.1
An $\varepsilon$-nilpotent structure $N$ of $(E,h)$ gives a nondegenerate section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$. Therefore $\xi_{N}$ in (2.4) does not depend on the choice of an admissible frame field $e$ of $N$, which was already obtained in [4].
Remark 2.2
Suppose $n=1$. The $2$-fold exterior power $\bigwedge^{2}\!E$ of $E$ is a vector bundle over $M$ of rank $6$ and $h$ induces a metric $\hat{h}$ of $\bigwedge^{2}\!E$ of signature (2,4). In addition, $\bigwedge^{2}\!E$ is decomposed as $\bigwedge^{2}\!E=\bigwedge^{2}_{+}\!E\oplus\bigwedge^{2}_{-}\!E$ by two subbundles +$\bigwedge^{2}_{+}\!E$, $\bigwedge^{2}_{-}\!E$ of rank $3$ and the restriction of $\hat{h}$ on each of them has signature (1,2). The light-like twistor spaces associated with $(E,h)$ are fiber bundles $U_{0}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ in ±$\bigwedge^{2}_{\pm}\!E$ respectively such that each fiber is a light cone. Each light-like line subbundle of +$\bigwedge^{2}_{+}\!E$ or $\bigwedge^{2}_{-}\!E$ corresponds to a light-like subbundle of $(E,h)$ of rank $2$ and each $\varepsilon$-nilpotent structure $N$ of $(E,h)$ corresponds to a section $\Omega_{N}$ of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ given by $(1/\sqrt{2})\xi_{N}$ ([3], [4]).
Remark 2.3
Suppose $n=1$. The space-like twistor spaces $U_{+}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ associated with $(E,h)$ are fiber bundles in ±$\bigwedge^{2}_{\pm}\!E$ respectively such that each fiber is a hyperboloid of two sheets. A section of $U_{+}\!\left(\textstyle\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponds to a complex structure of $E$ preserving $h$. See [2], [7] for the space-like twistor spaces. The time-like twistor spaces $U_{-}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ associated with $(E,h)$ are fiber bundles in ±$\bigwedge^{2}_{\pm}\!E$ respectively such that each fiber is a hyperboloid of one sheet. A section of $U_{-}\!\left(\textstyle\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponds to a paracomplex structure of $E$ reversing $h$. See [2], [22], [23] for the time-like twistor spaces. See [9], [16], [17] for the twistor spaces in the case $h$ is a Riemannian (i.e., positive-definite) metric, which are the prototypes of $U_{+}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$, $U_{-}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$ and $U_{0}\!\left(\textstyle\bigwedge^{2}_{\pm}\!E\right)$.
3 The Walker condition
Let $\nabla$ be an $h$-connection of $E$, i.e., a connection of $E$ satisfying $\nabla h=0$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. We say that $N$ satisfies the Walker condition with respect to $\nabla$ if for any local section $\psi$ of $\pi_{N}$, $\nabla\psi$ is a $1$-form valued in $\pi_{N}$. See [8], [11], [12], [15], [30] for Walker manifolds.
Let $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ be an admissible frame field of $N$. Let $\omega=[\omega^{i}_{j}]$ be the connection form of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$. Then we have $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. We represent $\omega$ as
$\omega=\left[\begin{array}[]{cccc}D_{11}&D_{12}&D_{13}&D_{14}\\ D_{21}&D_{22}&D_{23}&D_{24}\\ D_{31}&D_{32}&D_{33}&D_{34}\\ D_{41}&D_{42}&D_{43}&D_{44}\end{array}\right],$
where $D_{kl}$ is the $(k,l)$-block of $\omega$, which is an $n\times n$ matrix such that the components are given by
$\omega^{i}_{j}\quad(i=(k-1)n+1,\dots,kn,\ j=(l-1)n+1,\dots,ln).$
Since $\omega$ is a $1$-form valued in the Lie algebra of $SO(2n,2n)$, we have
$\begin{split}&{}^{t}\!D_{ii}=-D_{ii}\ (i=1,2,3,4),\quad{}^{t}\!D_{21}=-D_{12},\quad{}^{t}\!D_{43}=-D_{34},\\ &{}^{t}\!D_{31}=D_{13},\quad{}^{t}\!D_{41}=D_{14},\quad{}^{t}\!D_{32}=D_{23},\quad{}^{t}\!D_{42}=D_{24}.\end{split}$ (3.1)
In addition, $N$ satisfies the Walker condition with respect to $\nabla$ if and only if on a neighborhood of each point of $M$, an admissible frame field $e$ of $N$ satisfies
$\begin{split}D_{11}-D_{13}+D_{31}-D_{33}&=O_{n},\\ D_{21}-D_{23}-D_{41}+D_{43}&=O_{n},\\ D_{22}+D_{24}-D_{42}-D_{44}&=O_{n}.\end{split}$ (3.2)
The connection $\nabla$ induces connections of $\bigwedge^{2}\!E$ and $\bigwedge^{2n}\!E$, which are denoted by $\hat{\nabla}$. We will prove
Proposition 3.1
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then the following hold::::
  • (a)
    $N$ satisfies the Walker condition with respect to $\nabla$ if and only if ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ for a $1$-form $\alpha$ on $M;$
  • (b)
    $N$ is parallel with respect to $\nabla$ if and only if the corresponding section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$ is horizontal with respect to $\hat{\nabla}$, and if we suppose one of these conditions, then ^$\hat{\nabla}\xi_{N}=0$, and therefore $N$ satisfies the Walker condition with respect to $\nabla$.
Proof  If an $\varepsilon$-nilpotent structure $N$ of $(E,h)$ satisfies the Walker condition with respect to $\nabla$, then we have ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$, where $\alpha$ is a $1$-form on $M$ which is locally represented as
$\alpha:=\left(-\sum^{n}_{i=1}\omega^{2n+i}_{i}+\sum^{n}_{i=1}\omega^{3n+i}_{n+i}\right).$ (3.3)
If $N$ satisfies ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ for a $1$-form $\alpha$ on $M$, then each $\nabla\xi_{i}$ is a $1$-form valued in $\pi_{N}$, which means that $N$ satisfies the Walker condition. Hence we obtain (a) in Proposition 3.1. The condition $\nabla N=0$ is equivalent to $\omega\Lambda_{n}=\Lambda_{n}\omega$. This can be rewritten as
$\begin{split}&D_{11}-D_{13}=D_{22}-D_{42}=D_{33}-D_{31}=D_{44}-D_{24},\\ &D_{12}+D_{14}=-D_{21}+D_{41}=-D_{23}+D_{43}=-D_{32}-D_{34}.\end{split}$ (3.4)
We see that (3.4) is equivalent to ^$\hat{\nabla}\Theta_{N}=0$. Therefore $\nabla N=0$ is equivalent to ^$\hat{\nabla}\Theta_{N}=0$. If we suppose ^$\hat{\nabla}\Theta_{N}=0$, then by (2.4), we obtain ^$\hat{\nabla}\xi_{N}=0$. Hence we obtain (b) in Proposition 3.1. $\square$
Remark 3.1
It is already known that $\nabla N=0$ means that $N$ satisfies the Walker condition with respect to $\nabla$ (see [15]).
Remark 3.2
Suppose $n=1$. Then $\hat{\nabla}$ gives connections of ±$\bigwedge^{2}_{\pm}\!E$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then the corresponding section of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ is represented as $\Omega_{N}=(1/\sqrt{2})\xi_{N}$, and $\nabla N=0$ is equivalent to ^$\hat{\nabla}\xi_{N}=0$ ([3]). Since (3.1) means
$\omega^{i}_{i}=0\ (i=1,2,3,4),\quad\omega^{i+1}_{i}=-\omega^{i}_{i+1}\ (i=1,3),\quad\omega^{i}_{j}=\omega^{j}_{i}\ (|i-j|>1),$ (3.5)
the Walker condition is given by $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$ for an admissible frame field $e$ of $N$. In addition, noticing (3.3), we see that $N$ satisfies $\nabla N=0$ if and only if not only $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$ but also $\omega^{3}_{1}=\omega^{4}_{2}$ hold.
Remark 3.3
Suppose $n=1$. For a complex structure $I$ of $E$ preserving $h$, $\nabla I=0$ is equivalent to ^$\hat{\nabla}\Omega_{I}=0$ for the section $\Omega_{I}$ of $U_{+}\!\left(\textstyle\bigwedge^{2}_{+}\!E\right)$ or $U_{+}\!\left(\textstyle\bigwedge^{2}_{-}\!E\right)$ corresponding to $I$ ([2]). Similarly, for a paracomplex structure $J$ of $E$ reversing $h$, $\nabla J=0$ is equivalent to ^$\hat{\nabla}\Omega_{J}=0$ for the section $\Omega_{J}$ of $U_{-}\!\left(\textstyle\bigwedge^{2}_{+}\!E\right)$ or $U_{-}\!\left(\textstyle\bigwedge^{2}_{-}\!E\right)$ corresponding to $J$ ([2]).
Remark 3.4
Let $E$, $h$ be as in the beginning of Section 2. Let $J$ be a section of ${\rm End}\,E$. As in [4], we say that $J$ is an $\varepsilon$-paracomplex structure of $(E,h)$ if $J$ satisfies
  • (i)
    $J$ is a paracomplex structure of $E$,
  • (ii)
    $J$ is $h$-reversing, that is, $J^{*}h=-h$,
  • (iii)
    on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots e_{2n},e_{2n+1},\dots,e_{4n})$ of $E$ giving the orientation and satisfying
    $JeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\Lambda_{n,-},\quad\Lambda_{n,-}:=\left[\begin{array}[]{cccc}O_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\\ I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&-I_{n}&O_{n}&O_{n}\end{array}\right].$ (3.6)
Let $J$ be an $\varepsilon$-paracomplex structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $J$. We see that $J$ gives a section $J^{*}$ of ${\rm End}\,E^{*}$ by $J^{*}\phi^{*}=\phi^{*}\circ J$ for a local section $\phi^{*}$ of $E^{*}$. Then we have $J^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}\Lambda_{n,-}$. Therefore, if we set $\Theta_{J}(\phi^{*},\psi^{*}):=h^{*}(\phi^{*},J^{*}\psi^{*})$, then $\Theta_{J}$ is a section of $\bigwedge^{2}\!E$ and locally represented as
$\Theta_{J}=\sum^{n}_{i=1}(e_{i}\wedge e_{2n+i}+(\varepsilon 1)^{\delta_{i1}}e_{3n+i}\wedge e_{n+i}).$ (3.7)
Let $\nabla$ be an $h$-connection of $E$. Then $\nabla J=0$ is equivalent to ^$\hat{\nabla}\Theta_{J}=0$. Suppose that $\nabla J$ is locally represented as the tensor product of a $1$-form $\alpha$ and an $\varepsilon$-nilpotent structure $N$ so that $e$ is an admissible frame field of both $J$ and $N$. Then
$\begin{split}&\omega^{n+i}_{j}+(\varepsilon 1)^{\delta_{i1}}\omega^{3n+i}_{2n+j}=\omega^{n+i}_{2n+j}+(\varepsilon 1)^{\delta_{i1}}\omega^{3n+i}_{j}=\alpha\delta^{i}_{j},\\ &\omega^{i}_{j}-\omega^{2n+i}_{2n+j}=\omega^{i}_{2n+j}-\omega^{2n+i}_{j}=0,\\ &\omega^{n+i}_{n+j}-(\varepsilon 1)^{\delta_{i1}+\delta_{j1}}\omega^{3n+i}_{3n+j}=\omega^{n+i}_{3n+j}-(\varepsilon 1)^{\delta_{i1}+\delta_{j1}}\omega^{3n+i}_{n+j}=0\end{split}$ (3.8)
are obtained in [4] for $i,j=1,\dots,n$. Therefore by (3.7) and (3.8), we obtain ^$\hat{\nabla}\Theta_{J}=\alpha\otimes\Theta_{N}$.
Let $N$, $N^{\prime}$ be $\varepsilon$-nilpotent structures of $(E,h)$. Suppose that on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying
  • (i)
    $e$ is an admissible frame field of $N$,
  • (ii)
    $e^{\prime}:=(e_{1},\dots,e_{2n},-e_{2n+1},\dots,-e_{4n})$ is an admissible frame field of $N^{\prime}$.
Then $E=\pi_{N}\oplus\pi_{N^{\prime}}$. Referring to (3.2), we obtain
Proposition 3.2
Both of $N$ and $N^{\prime}$ satisfy the Walker condition if and only if the connection form of $\nabla$ with respect to $e$ as above satisfies not only (3.1) but also
$\begin{split}&D_{11}=D_{33},\quad D_{22}=D_{44},\quad D_{13}=D_{31},\quad D_{24}=D_{42},\\ &D_{43}=-D_{21},\quad D_{41}=-D_{23}.\end{split}$ (3.9)
Example 3.1
[…]
$F_{k}:=\left[\begin{array}[]{ccccc}a_{k}&b_{k}&0&\cdots&0\\ b_{k}&a_{k}&b_{k}&\ddots&\vdots\\ 0&b_{k}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&b_{k}\\ 0&\cdots&0&b_{k}&a_{k}\end{array}\right].$
Then we have $dF_{k}\wedge dF_{k}=O_{n}$ for $k=1,2$. We set
$\omega=\left[\begin{array}[]{cccc}O_{n}&O_{n}&dF_{1}&O_{n}\\ O_{n}&O_{n}&O_{n}&dF_{2}\\ dF_{1}&O_{n}&O_{n}&O_{n}\\ O_{n}&dF_{2}&O_{n}&O_{n}\end{array}\right].$
Then we have $d\omega+\omega\wedge\omega=O_{4n}$. Since we suppose that $\nabla$ is flat, there exists an ordered pseudo-orthonormal frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. Then $\omega$ is the connection form of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ and it satisfies (3.1) and (3.9). Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$, and $N^{\prime}$ an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e^{\prime}$ is an admissible frame field of $N^{\prime}$. Then $N$, $N^{\prime}$ satisfy the Walker condition. We have
^$\hat{\nabla}\xi_{N}=n(-da_{1}+da_{2})\otimes\xi_{N},\quad\hat{\nabla}\xi_{N^{\prime}}=n(da_{1}-da_{2})\otimes\xi_{N^{\prime}}.$
Therefore if $da_{1}\not=da_{2}$, then ^$\hat{\nabla}\xi_{N}$, ^$\hat{\nabla}\xi_{N^{\prime}}$ do not become zero and therefore by (b) in Proposition 3.1, $N$, $N^{\prime}$ are not parallel with respect to $\nabla$.
Example 3.2
Let $M$, $\nabla$ be as in Example 3.1. Let $a_{1}$, $a_{2}$ be functions on $M$ satisfying $da_{1}\not=da_{2}$. Suppose $n=1$ and set
$\omega=\left[\begin{array}[]{cccc}0&0&da_{1}&0\\ 0&0&0&da_{2}\\ da_{1}&0&0&0\\ 0&da_{2}&0&0\end{array}\right].$
Then we have $d\omega+\omega\wedge\omega=O_{4}$. Therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},e_{2},e_{3},e_{4})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4,\varepsilon}=eI^{\prime}_{4,\varepsilon}\omega$. Then $\omega$ is the connection form of $\nabla$ with respect to $eI^{\prime}_{4,\varepsilon}$ and it satisfies (3.1) and (3.9). Let $N$, $N^{\prime}$ be as in Example 3.1. Then $N$, $N^{\prime}$ satisfy the Walker condition. We have
^$\hat{\nabla}\xi_{N}=(-da_{1}+da_{2})\otimes\xi_{N},\quad\hat{\nabla}\xi_{N^{\prime}}=(da_{1}-da_{2})\otimes\xi_{N^{\prime}}.$
Therefore ^$\hat{\nabla}\xi_{N}$, ^$\hat{\nabla}\xi_{N^{\prime}}$ do not become zero and therefore $N$, $N^{\prime}$ are not parallel with respect to $\nabla$.
Example 3.3
Let $M$, $\nabla$ be as in Example 3.1 and suppose $n=1$. Let $N$ be an $\varepsilon$-nilpotent structure of $E$ with the Walker condition and $e$ an admissible frame field of $N$ defined on $M$. Then the connection form $\omega=[\omega^{i}_{j}]$ of $\nabla$ with respect to $eI^{\prime}_{4,\varepsilon}$ satisfies $d\omega+\omega\wedge\omega=0$ and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$. We can rewrite $d\omega+\omega\wedge\omega=0$ into
$\begin{split}d\omega^{2}_{1}+\omega^{3}_{2}\wedge\omega^{3}_{1}+\omega^{4}_{2}\wedge\omega^{4}_{1}&=0,\\ d\omega^{3}_{1}+\omega^{3}_{2}\wedge\omega^{2}_{1}-\omega^{4}_{3}\wedge\omega^{4}_{1}&=0,\\ d\omega^{4}_{1}+\omega^{4}_{2}\wedge\omega^{2}_{1}+\omega^{4}_{3}\wedge\omega^{3}_{1}&=0,\\ d\omega^{3}_{2}-\omega^{3}_{1}\wedge\omega^{2}_{1}-\omega^{4}_{3}\wedge\omega^{4}_{2}&=0,\\ d\omega^{4}_{2}-\omega^{4}_{1}\wedge\omega^{2}_{1}+\omega^{4}_{3}\wedge\omega^{3}_{2}&=0,\\ d\omega^{4}_{3}+\omega^{4}_{1}\wedge\omega^{3}_{1}+\omega^{4}_{2}\wedge\omega^{3}_{2}&=0.\end{split}$ (3.10)
Using the second and fifth equations in (3.10) and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$, we find a function $a$ satisfying
$da=\omega^{3}_{1}-\omega^{4}_{2};$ (3.11)
using the first and sixth equations in (3.10) and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$, we obtain
$d(\omega^{2}_{1}+\omega^{4}_{3})+(\omega^{2}_{1}+\omega^{4}_{3})\wedge da=0$ (3.12)
(using the third and fourth equations in (3.10) and $\omega^{1}_{2}+\omega^{1}_{4}+\omega^{3}_{2}+\omega^{3}_{4}=0$, we can obtain (3.12) again). From (3.12), we see that there exists a function $b$ satisfying
$\omega^{2}_{1}+\omega^{4}_{3}=\omega^{4}_{1}+\omega^{3}_{2}=e^{a}db.$ (3.13)
From (3.11) and (3.13), we obtain
$\omega^{3}_{2}=e^{a}db-\omega^{4}_{1},\quad\omega^{4}_{3}=e^{a}db-\omega^{2}_{1},\quad\omega^{4}_{2}=\omega^{3}_{1}-da.$ (3.14)
Applying (3.14) to (3.10), we obtain
$\begin{split}&d\omega^{2}_{1}+2\omega^{3}_{1}\wedge\omega^{4}_{1}+e^{a}db\wedge\omega^{3}_{1}-da\wedge\omega^{4}_{1}=0,\\ &d\omega^{3}_{1}+2\omega^{2}_{1}\wedge\omega^{4}_{1}+e^{a}db\wedge(\omega^{2}_{1}-\omega^{4}_{1})=0,\\ &d\omega^{4}_{1}-2\omega^{2}_{1}\wedge\omega^{3}_{1}-da\wedge\omega^{2}_{1}+e^{a}db\wedge\omega^{3}_{1}=0.\end{split}$ (3.15)
Therefore the connection form $\omega=[\omega^{i}_{j}]$ satisfies (3.15) for functions $a$, $b$ on $M$. Conversely, if $1$-forms $\omega^{2}_{1}$, $\omega^{3}_{1}$, $\omega^{4}_{1}$ and functions $a$, $b$ on $M$ satisfy (3.15), then $\omega=[\omega^{i}_{j}]$ defined by (3.5) and (3.14) satisfies $d\omega+\omega\wedge\omega=0$ and therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},e_{2},e_{3},e_{4})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4,\varepsilon}=eI^{\prime}_{4,\varepsilon}\omega$. Then $e$ defines an $\varepsilon$-nilpotent structure $N$ of $E$ such that $e$ is an admissible frame field of $N$. Since $\omega^{i}_{j}$, $a$, $b$ satisfy (3.13), $N$ satisfies the Walker condition. In particular,
  • (a)
    $a$ is constant if and only if $\nabla N=0$, that is, ^$\hat{\nabla}\xi_{N}=0$;
  • (b)
    $b$ is constant if and only if there exists an $\varepsilon$-nilpotent structure $N^{\prime}$ of $E$ satisfying the Walker condition such that $e^{\prime}$ is an admissible frame field of $N^{\prime}$.
Suppose that $b$ is constant. Then (3.15) is represented as
$\begin{split}&d\omega^{2}_{1}+2\omega^{3}_{1}\wedge\omega^{4}_{1}-da\wedge\omega^{4}_{1}=0,\\ &d\omega^{3}_{1}+2\omega^{2}_{1}\wedge\omega^{4}_{1}=0,\\ &d\omega^{4}_{1}-2\omega^{2}_{1}\wedge\omega^{3}_{1}-da\wedge\omega^{2}_{1}=0.\end{split}$ (3.16)
If $\omega^{3}_{1}=dc$ for a function $c$ on $M$, then
$\omega^{2}_{1}=\dfrac{1}{2}(e^{a-2c}d\phi+e^{-a+2c}d\psi),\quad\omega^{4}_{1}=\dfrac{1}{2}(e^{a-2c}d\phi-e^{-a+2c}d\psi)$ (3.17)
for functions $\phi$, $\psi$ on $M$ satisfying $d\phi\wedge d\psi=0$. Conversely, for functions $c$, $\phi$, $\psi$ on $M$ with $d\phi\wedge d\psi=0$, $1$-forms $\omega^{2}_{1}$, $\omega^{4}_{1}$ defined as in (3.17) and $\omega^{3}_{1}:=dc$ satisfy (3.16) (see Example 3.2 for the case where $\phi=\psi=0$ and $da\not=0$).
Example 3.4
Suppose $n\geq 2$. Let $M$, $\nabla$ be as in Example 3.1. Let $a_{1},\dots,a_{n},b_{1},\dots,b_{n}$ be functions on $M$ satisfying
$\sum^{n}_{i=1}a_{i}=\sum^{n}_{i=1}b_{i}.$ (3.18)
We set
$A:=\left[\begin{array}[]{ccccc}a_{1}&0&\cdots&0\\ 0&a_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&a_{n}\end{array}\right],\quad B:=\left[\begin{array}[]{ccccc}b_{1}&0&\cdots&0\\ 0&b_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&b_{n}\end{array}\right]$
and
$\omega=\left[\begin{array}[]{cccc}O_{n}&O_{n}&dA&O_{n}\\ O_{n}&O_{n}&O_{n}&dB\\ dA&O_{n}&O_{n}&O_{n}\\ O_{n}&dB&O_{n}&O_{n}\end{array}\right].$
Then we have $d\omega+\omega\wedge\omega=O_{4n}$. Therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$, and $N^{\prime}$ an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e^{\prime}$ is an admissible frame field of $N^{\prime}$. Then $N$, $N^{\prime}$ satisfy the Walker condition, by (3.9). By (3.3) and (3.18), we have ^$\hat{\nabla}\xi_{N}=0$ and ^$\hat{\nabla}\xi_{N^{\prime}}=0$. If $dA\not=dB$, then by (3.4), $N$, $N^{\prime}$ are not parallel with respect to $\nabla$.
Example 3.5
Let $M$, $\nabla$ be as in Example 3.1. Let $F$ be an $n\times n$ symmetric matrix such that each component is a function on $M$ and suppose $dF\not=O_{n}$ and $dF\wedge dF=O_{n}$. We set
$\omega=\left[\begin{array}[]{cccc}O_{n}&O_{n}&O_{n}&dF\\ O_{n}&O_{n}&dF&O_{n}\\ O_{n}&dF&O_{n}&O_{n}\\ dF&O_{n}&O_{n}&O_{n}\end{array}\right].$ (3.19)
Then we have $d\omega+\omega\wedge\omega=O_{4n}$. Therefore there exists an ordered pseudo-orthonormal frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ and satisfying $\nabla eI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\omega$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$. Then $N$ does not satisfy the Walker condition, because $\omega$ in (3.19) does not satisfy (3.2). We have
$(\nabla N)eI^{\prime}_{4n,\varepsilon}=2eI^{\prime}_{4n,\varepsilon}\left[\begin{array}[]{cccc}O_{n}&O_{n}&dF&O_{n}\\ O_{n}&O_{n}&O_{n}&-dF\\ dF&O_{n}&O_{n}&O_{n}\\ O_{n}&-dF&O_{n}&O_{n}\end{array}\right].$
In particular, if $dF=(1/2)dfI_{n}$ for a function $f$, then $\nabla N=df\otimes J$, where $J$ is an $\varepsilon$-paracomplex structure of $(E,h)$ defined by (3.6), and ^$\hat{\nabla}\Theta_{N}=df\otimes\Theta_{J}$. Therefore, if we suppose $n=1$, then ^$\hat{\nabla}\Omega_{N}=df\otimes\Omega_{J}$, where $\Omega_{N}$ is the section of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $N$ and given by $(1/\sqrt{2})\xi_{N}$, and $\Omega_{J}$ is the section of $U_{-}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $J$.
4 Special nilpotent structures
Let $E$ be an oriented vector bundle over $M$ of rank $4n$ and $h$ a neutral metric of $E$.
Definition 4.1
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then for a Lie subgroup $K$ of $SO(2n,2n)$, $N$ is called a $K$-nilpotent structure of $(E,h)$ if there exist an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of admissible frame fields of $N$ satisfying
  • (i)
    each $e_{\lambda}$ is defined on $U_{\lambda}$,
  • (ii)
    if $U_{\lambda}\cap U_{\mu}\not=\emptyset$ for $\lambda$, $\mu\in\Lambda$, then $e_{\mu}I^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$ for a function $A_{\lambda\mu}$ valued in $K$.
Let $N$ be a $K$-nilpotent structure of $(E,h)$. By definition, $N$ gives a principal $K$-bundle $P$ over $M$ associated with $E$ such that each $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is a local section of $P$ (see [26] for principal fiber bundles). Let $e$ be an admissible frame field of $N$ defined on an open set $U$ of $M$. Then $e$ is called a $K$-admissible frame field if $eI^{\prime}_{4n,\varepsilon}$ gives a local section of $P$, that is, if for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, there exists a $K$-valued function $A_{\lambda}$ on $U\cap U_{\lambda}$ satisfying $eI^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda}$. Based on Definition 4.1, we obtain
Proposition 4.1
Any $\varepsilon$-nilpotent structure of $(E,h)$ is a $G$-nilpotent structure of $(E,h)$.
Let $\nabla$ be an $h$-connection of $E$.
Definition 4.2
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then for a Lie subgroup $K$ of $SO(2n,2n)$, $N$ is called a $K$-nilpotent structure of $(E,h,\nabla)$ if there exist an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 and
  • (iii)
    the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $K$.
Let $N$ be a $K$-nilpotent structure of $(E,h,\nabla)$. Then $\nabla$ gives a connection in the principal $K$-bundle $P$. Let $e$ be a $K$-admissible frame field of $N$. Then the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ is given by $\omega=A^{-1}_{\lambda}\omega_{\lambda}A_{\lambda}+A^{-1}_{\lambda}dA_{\lambda}$. Therefore $\omega$ is valued in the Lie algebra of $K$.
Suppose that $\nabla$ is flat, that is, the curvature tensor of $\nabla$ vanishes. Then there exists an ordered pseudo-orthonormal local frame field $e$ of $(E,h)$ on a neighborhood $U$ of each point of $M$ giving the orientation of $E$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$. Then the connection form of $\nabla$ with respect to $e$ vanishes. Therefore for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, if we define an $SO(2n,2n)$-valued function $\tilde{A}_{\lambda}$ on $U\cap U_{\lambda}$ by $eI^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}\tilde{A}_{\lambda}$, then $\tilde{A}_{\lambda}$ satisfies $d\tilde{A}_{\lambda}=-\omega_{\lambda}\tilde{A}_{\lambda}$. This relation means that $d\tilde{A}_{\lambda}$ is valued in the distribution $\mathscr{D}$ on $SO(2n,2n)$ obtained by the right translations of the Lie algebra of $K$. If we suppose that $U\cap U_{\lambda}$ is connected, then there exists an integral manifold of $\mathscr{D}$ containing $\tilde{A}_{\lambda}$. We can choose $e$ so that $\tilde{A}_{\lambda}$ is valued in $K$ at a point. Since $K$ is an integral manifold of $\mathscr{D}$, $\tilde{A}_{\lambda}$ is valued in $K$ over its domain. If we suppose $K\subset G$, then $e$ is a $K$-admissible frame field of $N$. Hence we obtain
Proposition 4.2
Suppose that $\nabla$ is flat. Let $K$ be a Lie subgroup of $G$. Then for a $K$-nilpotent structure $N$ of $(E,h,\nabla)$, there exists a $K$-admissible frame field $e$ of $N$ on a neighborhood of each point of $M$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$.
Remark 4.1
Let $\eta_{1},\dots,\eta_{n},\eta_{n+1},\dots,\eta_{2n}$ be $2n$ vectors of $E^{4n}_{2n}$ defined by
$\Lambda_{n}=[\eta_{n+1}\ \dots\ \eta_{2n}\ -\eta_{1}\ \dots\ -\eta_{n}\ \eta_{n+1}\ \dots\ \eta_{2n}\ \eta_{1}\ \dots\ \eta_{n}].$
Let $W$ be a $2n$-dimensional subspace of $E^{4n}_{2n}$ spanned by $\eta_{1},\dots,\eta_{n},\eta_{n+1},\dots,\eta_{2n}$. Then $W$ is a light-like subspace of $E^{4n}_{2n}$. Let $SO(2n,2n)_{W}$ be a Lie subgroup of $SO(2n,2n)$ which consists of elements preserving $W$. Then we have $G\subset SO(2n,2n)_{W}$ and $G\not=SO(2n,2n)_{W}$ ([4]). An $SO(2n,2n)_{W}$-nilpotent structure of $(E,h,\nabla)$ is just an $\varepsilon$-nilpotent structure of $(E,h)$ satisfying the Walker condition. If $K=SO(2n,2n)_{W}$, then the conclusion of Proposition 4.2 is not necessarily valid, because of Example 3.1 $\sim$ Example 3.4, while for an almost $\varepsilon$-nilpotent structure $N$ of $E^{4n}_{2n}$ satisfying the Walker condition with respect to the Levi-Civita connection, $\pi_{N}$ gives a constant $2n$-dimensional light-like subspace of $E^{4n}_{2n}$.
We will prove
Proposition 4.3
A $G$-nilpotent structure of $(E,h,\nabla)$ is just an $\varepsilon$-nilpotent structure of $(E,h)$ parallel with respect to $\nabla$.
Proof  Let $N$ be a $G$-nilpotent structure of $(E,h,\nabla)$. Then the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ for any admissible frame field $e$ of $N$ is valued in the Lie algebra of $G$. Therefore we have $\omega\Lambda_{n}=\Lambda_{n}\omega$. By this, we obtain $\nabla(Ne)=N\nabla e$. Therefore $N$ is parallel with respect to $\nabla$. Suppose that $N$ is an $\varepsilon$-nilpotent structure of $(E,h)$ parallel with respect to $\nabla$. Then we have $\omega\Lambda_{n}=\Lambda_{n}\omega$ for the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ for any admissible frame field $e$ of $N$. This means that $\omega$ is valued in the Lie algebra of $G$. Therefore we can find an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 and (iii) in Definition 4.2 for $K=G$. Therefore $N$ is a $G$-nilpotent structure of $(E,h,\nabla)$. $\square$
Let $H$ be a Lie subgroup of $SO(2n,2n)$ defined by
$H:=\left\{\left[\begin{array}[]{cccc}A_{11}&-A_{21}&A_{31}&A_{41}\\ A_{21}&A_{11}&-A_{41}&A_{31}\\ A_{31}&-A_{41}&A_{11}&A_{21}\\ A_{41}&A_{31}&-A_{21}&A_{11}\end{array}\right]\in SO(2n,2n)\right\}.$ (4.1)
Then $H$ is a Lie subgroup of $G$. If $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then for an $H$-admissible frame field $e$ of $N$, the connection form $\omega$ of $\nabla$ with respect to $eI^{\prime}_{4n,\varepsilon}$ is represented as
$\omega=\left[\begin{array}[]{cccc}D_{11}&-D_{21}&D_{31}&D_{41}\\ D_{21}&D_{11}&-D_{41}&D_{31}\\ D_{31}&-D_{41}&D_{11}&D_{21}\\ D_{41}&D_{31}&-D_{21}&D_{11}\end{array}\right]\qquad\left(\begin{array}[]{l}{}^{t}D_{11}=-D_{11},\\ {}^{t}D_{21}=D_{21},\\ {}^{t}D_{31}=D_{31},\\ {}^{t}D_{41}=D_{41}\end{array}\right).$ (4.2)
The main objects of study in this section are $H$-nilpotent structures of $(E,h,\nabla)$.
Let $I$ be a complex structure of $E$ and $J_{1}$, $J_{2}$ paracomplex structures of $E$. Suppose that $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$, that is, suppose that $I$, $J_{1}$, $J_{2}$ satisfy
  • (i)
    $IJ_{1}=J_{2}$
and span a subbundle $E_{I,J_{1},J_{2}}$ of ${\rm End}\,E$ of rank $3$. In addition, suppose that $h$ is adapted to $E_{I,J_{1},J_{2}}$, that is, suppose
  • (ii)
    $I$ preserves $h$ and $J_{1}$, $J_{2}$ reverse $h$, that is, $h$ is Hermitian with respect to $I$ and paraHermitian with respect to $J_{1}$, $J_{2}$.
See [7], [21] for paraquaternionic structures (notice the sign in (i)). We say that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$ if $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ satisfy (i), (ii) and
  • (iii)
    $I$, $J_{1}$, $J_{2}$ are parallel with respect to $\nabla$.
See [13], [25] for neutral hyperKähler $4$-manifolds.
We will prove
Theorem 4.4
The following hold::::
  • (a)
    the section of $E_{I,J_{1},J_{2}}$ given by $N_{r,\theta}:=r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ for $r\in\mbox{\boldmath{$R$}}\setminus\{0\}$ and $\theta\in[0,2\pi)$ is an $H$-nilpotent structure of $(E,h);$
  • (b)
    if $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$, then $N_{r,\theta}$ is an $H$-nilpotent structure of $(E,h,\nabla)$.
Proof  Let $\{U_{\lambda}\}_{\lambda\in\Lambda}$ be an open cover of $M$ such that on each $U_{\lambda}$, there exists an ordered pseudo-orthonormal local frame field $e_{\lambda}=(e_{\lambda,1},\dots,e_{\lambda,2n},e_{\lambda,2n+1},\dots,e_{\lambda,4n})$ of $E$ satisfying
$Ie_{\lambda,i}=e_{\lambda,n+i},\quad J_{1}e_{\lambda,i}=e_{\lambda,2n+i},\quad J_{2}e_{\lambda,i}=e_{\lambda,3n+i}\quad(i=1,\dots,n).$ (4.3)
If $U_{\lambda}\cap U_{\mu}\not=\emptyset$ for $\lambda$, $\mu\in\Lambda$ and if we represent $e_{\mu}$ as $e_{\mu}=e_{\lambda}A_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$, then $A_{\lambda\mu}$ is valued in $H$. Hence $I$, $J_{1}$, $J_{2}$ define a principal $H$-bundle $P$ over $M$ associated with $E$ so that each $e_{\lambda}$ is a local section of $P$. For each $\lambda\in\Lambda$, let $N_{\lambda}$ be a section of the restriction of ${\rm End}\,E$ on $U_{\lambda}$ defined by $N_{\lambda}e_{\lambda}=e_{\lambda}\Lambda_{n}$. Then noticing that $H$ is contained in $G$, we can define a section $N$ of ${\rm End}\,E$ by $N=N_{\lambda}$ on each $U_{\lambda}$, which is an $\varepsilon$-nilpotent structure of $(E,h)$ for $\varepsilon=+$ or -- according to whether $e_{\lambda}$ gives the orientation of $E$ or not. In addition, $N$ is an $H$-nilpotent structure of $(E,h)$ and for each $\lambda\in\Lambda$, $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is an $H$-admissible frame field of $N$. By (4.3), $N$ coincides with $I+J_{2}=N_{1,0}$. For $\theta\in[0,2\pi)$, we set
$J^{\prime}_{1}:=(\cos\theta)J_{1}+(\sin\theta)J_{2},\quad J^{\prime}_{2}:=-(\sin\theta)J_{1}+(\cos\theta)J_{2}.$
Then $J^{\prime}_{1}$, $J^{\prime}_{2}$ are paracomplex structures of $E$ reversing $h$, and $I$, $J^{\prime}_{1}$, $J^{\prime}_{2}$ satisfy $IJ^{\prime}_{1}=J^{\prime}_{2}$ and span $E_{I,J_{1},J_{2}}$. Therefore referring to the above discussion, we obtain an $H$-nilpotent structure of $(E,h)$, which is given by
$I+J^{\prime}_{2}=I-(\sin\theta)J_{1}+(\cos\theta)J_{2}=N_{1,\theta}.$
For $t\not=0$ and $\delta\in\{1,-1\}$, we set
$I^{\prime}:=\delta(({\rm cosh}\,t)I+({\rm sinh}\,t)J^{\prime}_{2}),\quad J^{\prime\prime}_{2}:=\delta(({\rm sinh}\,t)I+({\rm cosh}\,t)J^{\prime}_{2}).$
Then $I^{\prime}$ is a complex structure of $E$ preserving $h$ and $J^{\prime\prime}_{2}$ is a paracomplex structure of $E$ reversing $h$. In addition, $I^{\prime}$, $J^{\prime}_{1}$ and $J^{\prime\prime}_{2}$ satisfy $I^{\prime}J^{\prime}_{1}=J^{\prime\prime}_{2}$ and span $E_{I,J_{1},J_{2}}$. Therefore referring to the above discussion, we obtain an $H$-nilpotent structure of $(E,h)$ given by
$I^{\prime}+J^{\prime\prime}_{2}=\delta e^{t}(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})=N_{\delta e^{t},\theta}.$
Hence we obtain (a) in Theorem 4.4. We set
$\begin{split}(\nabla e_{\lambda,1}\ \dots\ \nabla e_{\lambda,n})=&(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{11}+(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{21}\\ &+(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{31}+(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{41}.\end{split}$
Suppose that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$. Then by $\nabla I=0$, we obtain
$\begin{split}(\nabla e_{\lambda,n+1}\ \dots\ \nabla e_{\lambda,2n})=&-(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{21}+(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{11}\\ &-(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{41}+(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{31}.\end{split}$
Similarly, by $\nabla J_{1}=0$, we obtain
$\begin{split}(\nabla e_{\lambda,2n+1}\ \dots\ \nabla e_{\lambda,3n})=&(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{31}-(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{41}\\ &+(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{11}-(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{21},\end{split}$
and by $\nabla J_{2}=0$, we obtain
$\begin{split}(\nabla e_{\lambda,3n+1}\ \dots\ \nabla e_{\lambda,4n})=&(e_{\lambda,1}\ \dots\ e_{\lambda,n})D_{41}+(e_{\lambda,n+1}\ \dots\ e_{\lambda,2n})D_{31}\\ &+(e_{\lambda,2n+1}\ \dots\ e_{\lambda,3n})D_{21}+(e_{\lambda,3n+1}\ \dots\ e_{\lambda,4n})D_{11}.\end{split}$
Therefore the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}$ is given as in (4.2) and therefore the connection form $\omega_{\lambda}$ is valued in the Lie algebra of $H$. Hence $\nabla$ gives a connection in $P$ and $N=N_{1,0}$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Similarly, for $r\in\mbox{\boldmath{$R$}}\setminus\{0\}$ and $\theta\in[0,2\pi)$, $N_{r,\theta}$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Hence we obtain (b) in Theorem 4.4. $\square$
Remark 4.2
Whether $e_{\lambda}$ gives the orientation of $E$ does not depend on $\lambda\in\Lambda$, and $N_{r,\theta}$ is a $+$ or ---nilpotent structure according to whether $e_{\lambda}$ gives the orientation or not, as was already seen in the above proof.
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$ and suppose that $N$ is an $H$-nilpotent structure of $(E,h)$. Then the light-like subbundle $\pi_{N}$ of rank $2n$ is determined by $N$ and locally spanned by $\xi_{1},\dots,\xi_{2n}$ as in (2.3) for an $H$-admissible frame field $e$ of $N$. In addition, we will prove
Theorem 4.5
Let $N$ be as above.
  • (a)
    There exists a unique light-like subbundle $\pi^{\times}_{N}$ of $(E,h)$ of rank $2n$ which is locally spanned by
    $\begin{array}[]{lcl}\xi^{\prime}_{1}:=e_{1}+e_{2n+1},&&\xi^{\prime}_{i}:=e_{i}+e_{2n+i},\\ \xi^{\prime}_{n+1}:=e_{n+1}-\varepsilon e_{3n+1},&&\xi^{\prime}_{n+i}:=e_{n+i}-e_{3n+i}\end{array}\ (i=2,\dots,n)$ (4.4)
    for any $H$-admissible frame field $e$ of $N$. In particular, $E=\pi_{N}\oplus\pi^{\times}_{N}$.
  • (b)
    There exists a unique $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{\times}$ of $(E,h)$ such that each $H$-admissible frame field of $N^{\times}$ is given by $e^{\prime}=(e_{1},\dots,e_{2n},-e_{2n+1},\dots,-e_{4n})$ for an $H$-admissible frame field $e$ of $N$. In particular, $\pi_{N^{\times}}=\pi^{\times}_{N}$.
  • (c)
    The condition that $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$ is equivalent to the condition that $N^{\times}$ is an $H$-nilpotent structure of $(E,h,\nabla)$.
Proof  For an $H$-admissible frame field $e$ of $N$, the subbundle of the restriction of $E$ on the domain of $e$ spanned by $\xi^{\prime}_{i}$, $\xi^{\prime}_{n+i}$ ($i=1,\dots,n$) as in (4.4) does not depend on the choice of $e$, because of the definition of $H$ in (4.1). Therefore there exists a unique light-like subbundle $\pi^{\times}_{N}$ of $(E,h)$ of rank $2n$ satisfying (a) in Theorem 4.5. Hence we obtain (a) in Theorem 4.5. For $N$, we can find an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of $H$-admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 for $K=H$. We set
$A^{\prime}_{\lambda\mu}:=I_{2n,2n}A_{\lambda\mu}I_{2n,2n},\quad I_{2n,2n}:=\left[\begin{array}[]{cccc}I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&I_{n}&O_{n}&O_{n}\\ O_{n}&O_{n}&-I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\end{array}\right]$
for $\lambda$, $\mu\in\Lambda$ satisfying $U_{\lambda}\cap U_{\mu}\not=\emptyset$. Then $A^{\prime}_{\lambda\mu}$ satisfies $e^{\prime}_{\mu}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$, and since $A_{\lambda\mu}$ is valued in $H$, $A^{\prime}_{\lambda\mu}$ is also valued in $H$. Therefore there exists an $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{\times}$ of $(E,h)$ such that $e^{\prime}_{\lambda}$ ($\lambda\in\Lambda$) are $H$-admissible frame fields of $N^{\times}$ and then $N^{\times}$ is uniquely determined by $\{e^{\prime}_{\lambda}\}_{\lambda\in\Lambda}$. Let $\tilde{e}$ be an $H$-admissible frame field of $N^{\times}$ defined on an open set $U$ of $M$. Then for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, there exists an $H$-valued function $A^{\prime}_{\lambda}$ on $U\cap U_{\lambda}$ satisfying $\tilde{e}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda}$. This relation is equivalent to $\tilde{e}^{\prime}I^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda}$ with $A_{\lambda}:=I_{2n,2n}A^{\prime}_{\lambda}I_{2n,2n}$. Since $A_{\lambda}$ is $H$-valued, $\tilde{e}^{\prime}$ is an $H$-admissible frame field of $N$. Hence we obtain (b) in Theorem 4.5. Suppose that $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Then we can suppose that the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. Therefore the connection form $\omega^{\prime}_{\lambda}$ of $\nabla$ with respect to $e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. This means that $N^{\times}$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Similarly, we can prove the converse. Hence we obtain (c) in Theorem 4.5. $\square$
Let $N$ be an $H$-nilpotent structure of $(E,h)$. Then we call $N^{\times}$ as in Theorem 4.5 the dual $H$-nilpotent structure of $N$. We see that $N$ is the dual $H$-nilpotent structure of $N^{\times}$: $(N^{\times})^{\times}=N$.
For an $H$-nilpotent structure $N$ of $(E,h)$, we set
$I:=\dfrac{1}{2}(N+N^{\times}),\quad J_{2}:=\dfrac{1}{2}(N-N^{\times}),\quad J_{1}:=-IJ_{2}.$ (4.5)
Then $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$. In addition, if $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$. Hence we obtain
Corollary 4.6
An $H$-nilpotent structure $N$ of $(E,h)$ defines a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$ by (4.5). In addition, if $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$.
Remark 4.3
Let $E$, $h$ be as in the beginning of Section 2. Let $I$ be a section of ${\rm End}\,E$. We say that $I$ is an $\varepsilon$-complex structure of $(E,h)$ if $I$ satisfies
  • (i)
    $I$ is a complex structure of $E$,
  • (ii)
    $I$ is $h$-preserving, that is, $I^{*}h=h$,
  • (iii)
    on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots e_{2n},e_{2n+1},\dots,e_{4n})$ of $E$ giving the orientation and satisfying
    $IeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\Lambda_{n,+},\quad\Lambda_{n,+}:=\left[\begin{array}[]{cccc}O_{n}&-I_{n}&O_{n}&O_{n}\\ I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\\ O_{n}&O_{n}&I_{n}&O_{n}\end{array}\right].$
Let $I$ be an $\varepsilon$-complex structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $I$. We see that $I$ gives a section $I^{*}$ of ${\rm End}\,E^{*}$ by $I^{*}\phi^{*}=\phi^{*}\circ I$ for a local section $\phi^{*}$ of $E^{*}$. Then we have $I^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n,+}$. Therefore, if we set $\Theta_{I}(\phi^{*},\psi^{*}):=h^{*}(\phi^{*},I^{*}\psi^{*})$, then $\Theta_{I}$ is a section of $\bigwedge^{2}\!E$ and locally represented as
$\Theta_{I}=\sum^{n}_{i=1}(e_{i}\wedge e_{n+i}-(\varepsilon 1)^{\delta_{i1}}e_{2n+i}\wedge e_{3n+i}).$ (4.6)
For an $h$-connection $\nabla$ of $E$, $\nabla I=0$ is equivalent to ^$\hat{\nabla}\Theta_{I}=0$. Let $N$ be an $\varepsilon$-nilpotent and $H$-nilpotent structure of $(E,h)$ and $e$ an $H$-admissible frame field of $N$. Let $I$, $J_{2}$ be as in (4.5). Then $I$ is an $\varepsilon$-complex structure of $(E,h)$ such that $e$ is an admissible frame field of $I$ and $J_{2}$ is an $\varepsilon$-paracomplex structure of $(E,h)$ in the sense of Remark 3.4 such that
$(e_{n+1},\dots,e_{2n},-e_{1},\dots,-e_{n},e_{2n+1},\dots,e_{3n},e_{3n+1},\dots,e_{4n})$
is an admissible frame field of $J_{2}$. In addition, $\Theta_{N}=\Theta_{I}+\Theta_{J_{2}}$. Let $J_{1}$ be as in (4.5). Then $J_{1}$ is an $\varepsilon$-paracomplex structure of $(E,h)$ in the sense of Remark 3.4 such that $e$ is an admissible frame field of $J_{1}$.
Remark 4.4
Let $I$ be a complex structure of $E$ and $J_{1}$, $J_{2}$ paracomplex structures of $E$. Suppose that $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$. Then $N:=I+J_{2}$ is an $H$-nilpotent structure of $(E,h)$ and $N^{\times}:=I-J_{2}$ is its dual $H$-nilpotent structure.
Remark 4.5
If $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then we see by (4.2) that the connection form of an $H$-admissible frame field $e$ of $N$ satisfies (3.9).
Remark 4.6
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Let $e$, $\tilde{e}$ be admissible frame fields of $N$ and $A$ a $G$-valued function on the intersection of the domains of $e$, $\tilde{e}$ given by $\tilde{e}I^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}A$. Then the subbundle spanned by $\xi_{i}$, $\xi_{n+i}$ ($i=1,\dots,n$) coincides with the subbundle spanned by $\tilde{\xi}_{i}$, $\tilde{\xi}_{n+i}$ ($i=1,\dots,n$). In addition, if the subbundle spanned by $\xi^{\prime}_{i}$, $\xi^{\prime}_{n+i}$ ($i=1,\dots,n$) coincides with the subbundle spanned by $\tilde{\xi}^{\prime}_{i}$, $\tilde{\xi}^{\prime}_{n+i}$ ($i=1,\dots,n$), then $A$ is valued in $H$.